### v2

# Masked Language Modeling Becomes Conditional Density Estimation for Tabular Data Synthesis

Abstract
In this paper, our goal is to generate synthetic data for heterogeneous (mixed-type) tabular datasets with high machine learning utility (MLu). Since the MLu performance depends on accurately approximating the conditional distributions, we focus on devising a synthetic data generation method based on conditional distribution estimation. We introduce MaCoDE by redefining the consecutive multi-class classification task of Masked Language Modeling (MLM) as histogram-based non-parametric conditional density estimation. Our approach enables the estimation of conditional densities across arbitrary combinations of target and conditional variables. We bridge the theoretical gap between distributional learning and MLM by demonstrating that minimizing the orderless multi-class classification loss leads to minimizing the total variation distance between conditional distributions. To validate our proposed model, we evaluate its performance in synthetic data generation across 10 real-world datasets, demonstrating its ability to adjust data privacy levels easily without re-training. Additionally, since masked input tokens in MLM are analogous to missing data, we further assess its effectiveness in handling training datasets with missing values, including multiple imputations of the missing entries.
1 Introduction
There are two main objectives in synthetic tabular data generation: (1) preserving the statistical characteristics of the original dataset and (2) achieving comparable machine learning utility (MLu) to the original dataset. In this paper, our focus is on generating synthetic data with high MLu performance. Note that achieving high statistical fidelity does not guarantee high MLu performance (Hansen et al. 2023).
Given that MLu performance depends on accurately approximating conditional distributions, we focus on developing a synthetic data generation method based on conditional distribution estimation. However, two important properties of tabular data must be considered: ($i$) tabular data can consist of mixed types of data (Borisov et al. 2021; Shwartz-Ziv and Armon 2022), and ($ii$) the tabular data does not have an intrinsic ordering among columns (Gulati and Roysdon 2023).
($i$) Considering the heterogeneous nature of tabular data and aiming to develop a method that addresses the challenges of modeling diverse distributions of continuous columns, we employ histogram-based non-parametric conditional density estimation through a multi-class classification task (Li, Bondell, and Reich 2019). This approach enables us to apply the classification loss uniformly across all types of columns. Since the histogram-based approach is theoretically valid only when continuous variables have bounded supports (Wasserman 2006; Li, Bondell, and Reich 2019), we transform continuous columns using the Cumulative Distribution Function (CDF) and constrain their values to the interval $[0,1]$ (Li et al. 2021; Fang et al. 2022).
($ii$) To learn the arbitrary generation ordering of columns, we utilize the Masked Language Modeling (MLM) approach (Devlin et al. 2019). By employing a masking scheme and the BERT model architecture, our proposed model enables the estimation of conditional densities across arbitrary combinations of target and conditional variables (Ghazvininejad et al. 2019; Ivanov, Figurnov, and Vetrov 2019; Nazábal et al. 2020). (Gulati and Roysdon 2023) proposed a similar method called TabMT, however, TabMT faces challenges in distributional learning because it relies on predicting the K-means cluster index of masked entries. Our approach contrasts with existing auto-regressive density estimators, which generate data in a fixed column order (Hansen 1994; Kamthe, Assefa, and Deisenroth 2021; Letizia and Tonello 2022). Additionally, (Germain et al. 2015; Papamakarios, Pavlakou, and Murray 2017) are also able to estimate conditional densities but differ from our approach by masking the model weights rather than the input.
Therefore, our proposed method redefines the consecutive multi-class classification task of MLM as histogram-based non-parametric conditional density estimation. We term our proposed model MaCoDE (Masked Conditional Density Estimation). The main contribution of our work is bridging the theoretical gap between distributional learning and the consecutive minimization of the multi-class classification loss within the MLM approach.
Specifically, we demonstrate that minimizing the orderless multi-class classification loss, when combined with the CDF transformation, provides theoretical validity for minimizing the discrepancy between conditional distributions in terms of total variation distance. This implies that we do not need to consider the ordering of bins, which could otherwise serve as a useful inductive bias. Note that, in the natural language domain, previous attempts to interpret MLM as distributional learning have been somewhat limited, relying on pseudo-likelihood or Markov random fields (Ghazvininejad et al. 2019; Wang and Cho 2019; Salazar et al. 2019; Ng, Cho, and Ghassemi 2020; Hennigen and Kim 2023).
We substantiate the effectiveness of our proposed method by evaluating its performance in synthetic data generation across 10 real-world tabular datasets, demonstrating its capability to adjust data privacy levels easily without re-training. Given that masked input tokens in MLM can be viewed as missing data, we also assess our model’s effectiveness in handling training datasets with missing values. Moreover, as our proposed model estimates the conditional distribution while accommodating arbitrary conditioning sets, it can address various missingness patterns - an essential capability for generating samples and performing multiple imputations (van Buuren 2012; Ivanov, Figurnov, and Vetrov 2019; Nazábal et al. 2020). Consequently, we further validate our method’s effectiveness by evaluating its performance in multiple imputations across various missing data mechanisms.
2 Related Works
Existing methods using deep generative models aim to directly minimize the discrepancy between the multivariate ground-truth distribution and the generative model. These include CTGAN (Xu et al. 2019), TVAE (Xu et al. 2019), CTAB-GAN (Zhao et al. 2021), CTAB-GAN+ (Zhao et al. 2023b), DistVAE (An and Jeon 2023), and TabDDPM (Kotelnikov et al. 2023).
In the realm of transformer-based synthesizers, methods such as TabPFGen (Ma et al. 2023), TabMT (Gulati and Roysdon 2023), and REaLTabFormer (Solatorio and Dupriez 2023) have been proposed. TabMT utilizes an MLM-based approach to generate synthetic data by predicting cluster indices from K-means clustering. TabPFGen is an energy-based model that leverages the Bayesian inference of TabPFN (Müller et al. 2022) framework. REaLTabFormer employs an autoregressive Transformer architecture akin to GPT-2 (Radford et al. 2019), applying natural language generation techniques to tabular data. Additionally, methods such as (Kamthe, Assefa, and Deisenroth 2021; Letizia and Tonello 2022; Drouin, Marcotte, and Chapados 2022; Ashok et al. 2024) use transformers for copula density estimation, specifically in time-series datasets.
3 Proposal
Figure 1: Overall structure of MaCoDE. In this case, the value of the second column is masked (replaced with ‘0’) and predicted.
Notations. Let ${\bf x}\in\mathbb{R}^{p}$ denote an observation consisting of continuous and categorical (discrete) variables, and the $j$th variable (column) is denoted as ${\bf x}_{j}$. Here, subscript $j$ refers to the $j$th element. $I_{C}$ and $I_{D}$ represent the index sets for continuous and categorical variables, where $I_{C}\cup I_{D}=\{1,\cdots,p\}$. The observed dataset is denoted as $\{{\bf x}^{(i)}\}_{i=1}^{n}$. ${\bf m}\in\{0,1\}^{p}$ is a binary vector indicating masked values, with ${\bf m}_{j}=0$ indicating the $j$th column is masked (if ${\bf m}_{j}=1$, then the $j$th column is not masked). $F^{*}_{j}$ indicates the ground-truth CDF of the $j$th column, and $\hat{F}_{j}$ is an estimator of $F^{*}_{j}$.
Overview. Without loss of generality, we can consider an arbitrary conditional density function: for $j\in I_{C}$,
$p_{j}^{*}({\bf x}_{j}|{\bf x}_{1},\cdots,{\bf x}_{j-1}).$ (1)
Our primary objective is to estimate (1). However, there are two major challenges in estimating (1):
  1. 1.
    Modeling non-uniform distributions of continuous columns, ${\bf x}_{j}$.
  2. 2.
    Handling arbitrary combinations of conditional variables, ${\bf x}_{1},\cdots,{\bf x}_{j-1}$.
Assumption 1.
For all $j\in I_{C}$, there exists $\hat{F}_{j}:\mathbb{R}\mapsto[0,1]$ such that $\hat{F}_{j}$ is invertible and differentiable.
In this paper, we address these major challenges by unifying a histogram-based conditional density estimation and the MLM approach. And we assume that $\hat{F}_{j}$ satisfying Assumption 1 is given. For $j\in I_{C}$,
$p_{j}^{*}({\bf x}_{j}|{\bf x}_{-j})$ $=$ $c^{*}_{j}(\hat{F}_{j}({\bf x}_{j})|{\bf x}_{-j})\cdot\hat{p}_{j}({\bf x}_{j})$ (2)
by the change of variable in terms of $\hat{F}_{j}$, where $c^{*}_{j}$ is the conditional density of $\hat{F}_{j}({\bf x}_{j})$, ${\bf x}_{-j}\coloneqq({\bf x}_{1},\cdots,{\bf x}_{j-1})$, and $\hat{p}_{j}$ denotes the density of $\hat{F}_{j}$. In particular, we estimate $c^{*}_{j}$ in (2) using a histogram-based approach.
Classification Target (Discretization)
The discretization is essential since the MLM-based approach hinges on multi-class classification tasks. In this section, we will describe how to transform the columns of a tabular dataset into classification targets. We denote the classification target (i.e., label) of the $j$th column as ${\bf y}_{j}$.
Continuous column. Firstly, since the histogram-based approach requires continuous columns to have bounded supports, we transform them into random variables within the $[0,1]$ range using their marginal CDFs. Then, we partition the $[0,1]$ interval with $L+1$ cut-points, $b_{0},b_{1},\cdots,b_{L}$, where $10=b_{0}<b_{1}<b_{2}<\cdots<b_{L-1}<b_{L}=$ 1, resulting in $L$ bins. We define the classification target ${\bf y}_{j}$ based on the interval within which $\hat{F}_{j}({\bf x}_{j})$ falls, as follows:
${\bf y}_{j}\coloneqq\sum_{s=0}^{L-1}\mathbb{I}\Big{(}b_{s}\leq\hat{F}_{j}({\bf x}_{j})\Big{)},$
indicating that the classification target is the bin index, where $\mathbb{I}(\cdot)$ represents the indicator function, and ${\bf y}_{j}\in[L]\coloneqq\{1,2,\cdots,L\}$. Thus, if $b_{l-1}\leq\hat{F}_{j}({\bf x}_{j})<b_{l}$, then ${\bf y}_{j}=l$.
Our discretization procedure presents an advantage: since the lower and upper bounds are naturally defined as 0 and 1 within our approach, determining the number of bins becomes more intuitive. For example, finer bins near the boundaries of 0 and 1 can be employed to achieve more accurate tail density estimation.
Categorical column. For categorical variables, without transformation procedures like those for continuous columns, we define ${\bf y}_{j}\coloneqq{\bf x}_{j}$ for $j\in I_{D}$. To simplify notation, we denote the number of levels for all categorical variables as $L$. Allowing categorical variables with different numbers of categories is straightforward.
Definition 1.
The discretization function $g:\mathbb{R}^{p}\mapsto[L]^{p}$ is defined as
$g({\bf x};\hat{F})_{j}$$\coloneqq$ $\begin{cases}\sum_{s=0}^{L-1}\mathbb{I}(b_{s}\leq\hat{F}_{j}({\bf x}_{j})),&\text{if $j\in I_{C}$}\\ {\bf x}_{j},&\text{if $j\in I_{D}$}\end{cases},$
for $j=1,\cdots,p$, where $\hat{F}\coloneqq\{\hat{F}_{j}:j\in I_{C}\}$.
Finally, we transform the observations ${\bf x}$ into the discretized label ${\bf y}$ using the discretization function $g$ of Definition 1, as ${\bf y}_{j}=g({\bf x};\hat{F})_{j}$ for all $j$.
Masked Conditional Density Estimation
Target distribution. For all $j\in\{1,2,\cdots,p\}$, the ground-truth conditional probability of $g({\bf x};\hat{F})_{j}=l$ given the other observed variables, i.e., $\{{\bf x}_{k}:{\bf m}_{k}=1,k\neq j\}$, is defined as
$\Pr\left(g({\bf x};\hat{F})_{j}=l\Big{|}\{{\bf x}_{k}:{\bf m}_{k}=1,k\neq j\}\right)$
$\eqqcolon$ $\pi^{*}_{jl}\Big{(}\{{\bf x}_{k}:{\bf m}_{k}=1,k\neq j\}\Big{)},$
which also corresponds to the conditional probability of $\hat{F}_{j}({\bf x}_{j})$ in the $l$th bin, $[b_{l-1},b_{l})$.
Then, the ground-truth conditional distribution of $g({\bf x};\hat{F})_{j}$, which is our target distribution, is written accordingly:
$p_{j}^{*}\Big{(}g({\bf x};\hat{F})_{j}|\{{\bf x}_{k}:{\bf m}_{k}=1,k\neq j\}\Big{)}$ (3)
$=$ $\prod_{l=1}^{L}\pi^{*}_{jl}\Big{(}\{{\bf x}_{k}:{\bf m}_{k}=1,k\neq j\}\Big{)}^{\mathbb{I}(g({\bf x};\hat{F})_{j}=l)}.$
MaCoDE. For $j$ such that ${\bf m}_{j}=0$, we parameterize (3) as follows:
$p_{j}\left(g({\bf x};\hat{F})_{j}|g({\bf x};\hat{F})\odot{\bf m};\theta\right)$
$=$ $\prod_{l=1}^{L}\pi_{jl}\left(g({\bf x};\hat{F})\odot{\bf m};\theta\right)^{\mathbb{I}(g({\bf x};\hat{F})_{j}=l)},$
where $\odot$ denotes element-wise multiplication and $\sum_{l=1}^{L}\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m};\theta)=1$ for all $j$. In this paper, we use the empirical CDF for $\hat{F}_{j}$. Here, $\theta$ represents all the parameters of the transformer encoder-based classifier, allowing us to process inputs of arbitrary lengths and accommodate different combinations of observed and masked variables. Additionally, we denote $(\pi_{j1}(\cdot;\theta),\cdots,\pi_{jL}(\cdot;\theta))$ as $\pi_{j}(\cdot;\theta)$.
Then, the objective function for a single observation is defined as the negative log-likelihood of masked entries:
${\cal L}({\bf y},{\bf m};\theta)$$\coloneqq$ $-\sum_{j:{\bf m}_{j}=0}\log p_{j}({\bf y}_{j}|{\bf y}\odot{\bf m};\theta)$
$=$ $-\sum_{j:{\bf m}_{j}=0}\sum_{l=1}^{L}\mathbb{I}({\bf y}_{j}=l)\cdot\log\pi_{jl}({\bf y}\odot{\bf m};\theta),$
where the label ${\bf y}_{j}$ is defined as ${\bf y}_{j}=g({\bf x};\hat{F})_{j}$ for all $j$. Our objective function (3) estimates the conditional distribution of $\hat{F}_{j}({\bf x}_{j})$ in each bin, where its target distribution is (3). In other words, our objective is to approximate the true conditional probability $\pi^{*}_{jl}$ by our estimated probability $\pi_{jl}$.
Similar to MLM, the bin index $({\bf y}\odot{\bf m})_{j}$ serves as a ‘word’ index aligned with the $j$th column, encompassing its own vocabulary set of $L+1$ words (i.e., $\{0,1,2,\cdots,L\}$), including ‘0’ for the masked input. And $\pi_{jl}$ represents the probability that the model outputs the word index $l$ from the $j$th column.
For each $j$, the embedding layer preprocesses of $({\bf y}\odot{\bf m})_{j}$ through one-hot encoding, with each bin index being assigned a learnable embedding vector. Note that our approach to handling masked entries shares similarities with the existing zero imputation technique, where all masked entries are replaced with zeros (Ivanov, Figurnov, and Vetrov 2019; Mattei and Frellsen 2019; Nazábal et al. 2020; Ipsen, Mattei, and Frellsen 2021). However, in our proposed method, although all masked entries are replaced with the same bin index, ‘0’, each imputed zero value is embedded through distinct embedding vectors for each column.
Definition 2 (Mask distribution).
The distribution of mask vector ${\bf m}$ is defined as:
[…]
where $p(u)$ is the uniform distribution density, and ${\bf m}_{1},{\bf m}_{2},\cdots,{\bf m}_{p}$ follow conditionally independent Bernoulli distributions with probability $u$ given $u$.
Finally, we minimize the following objective function with respect to $\theta$:
$\min_{\theta}\quad\sum_{i=1}^{n}\mathbb{E}_{p({\bf m})}\big{[}{\cal L}({\bf y}^{(i)},{\bf m};\theta)\big{]},$ (5)
where the discretized training dataset is $\{{\bf y}^{(i)}\}_{i=1}^{n}=\{g({\bf x}^{(i)};\hat{F})\}_{i=1}^{n}$, and $\mathbb{E}_{p({\bf m})}$ is approximated by Monte-Carlo and ancestral sampling. The distribution $p({\bf m})$ can be defined by the user based on the specific problem requirements (Ivanov, Figurnov, and Vetrov 2019). As in BERT (Devlin et al. 2019), the task of our objective function (5) is to predict the original label for all masked inputs. The overall structure of MaCoDE is outlined in Figure 1.
Remark 1.
We want to emphasize that our objective function (5) does not imply an assumption of conditional independence among masked entries given the observed data. Instead, by ensuring that $p({\bf m})$ has full support over $\{0,1\}^{p}$, sampling ${\bf m}$ from $p({\bf m})$ allows us to learn conditional densities encompassing all possible combinations of conditioning sets and target variables (Ivanov, Figurnov, and Vetrov 2019; Gulati and Roysdon 2023). Furthermore, this training scheme remains invariant to the missing data scenario.
Algorithm 1 Synthetic Data Generation
Initialize: For all $j$, $\hat{{\bf y}}_{j}\leftarrow 0$ and $\hat{{\bf m}}_{j}\leftarrow 0$. (all columns are masked)
Output: A synthetic sample $\hat{{\bf x}}$.
1:  for $j=randperm\{1,2,\cdots,p\}$ do {tabular data lacks the inherent ordering between columns}
2:     $\hat{{\bf y}}_{j}\sim Cat\big{(}\pi_{j}(\hat{{\bf y}}\odot\hat{{\bf m}};\theta)/\tau\big{)}$choosing a histogram bin (index) according to $\pi_{j}$}
3:     $\hat{{\bf m}}_{j}\leftarrow 1$un-mask the $j$th column}
4:  end for
5:  for $j=1,2,\cdots,p$ do
6:     if $j\in I_{C}$ then
7:        $u\sim U(b_{\hat{{\bf y}}_{j}-1},b_{\hat{{\bf y}}_{j}})$uniform sampling a quantile level from the bin interval}
8:        $\hat{{\bf x}}_{j}\leftarrow\hat{F}^{-1}_{j}(u)$inverse transform sampling of $\hat{p}_{j}$}
9:     end if
10:     if $j\in I_{D}$ then
11:        $\hat{{\bf x}}_{j}\leftarrow\hat{{\bf y}}_{j}$
12:     end if
13:  end for
Synthetic data generation. Tabular data lacks the inherent ordering between columns, unlike natural language. Therefore, as outlined in Algorithm 1, MaCoDE randomly generates one column at a time, conditioned on masked subset sizes from $p$ to $1$, in descending order ($p\rightarrow p-1\rightarrow\cdots\rightarrow 2\rightarrow 1$).
Controllable privacy level. Adjusting the privacy level during synthetic data generation is crucial in tabular domain (Park et al. 2018). Similar to TabMT (Gulati and Roysdon 2023), we can also regulate the privacy level using a single hyper-parameter $\tau$ without the need for re-training, while other existing synthesizers have fixed trade-offs between synthetic data quality and privacy level (Xu et al. 2019; Zhao et al. 2023b; Kotelnikov et al. 2023) or require re-training (An and Jeon 2023).
By increasing the temperature parameter $\tau$ in $\hat{{\bf y}}_{j}\sim Cat\big{(}\pi_{j}(\hat{{\bf y}}\odot\hat{{\bf m}};\theta)/\tau\big{)}$ at line 2 of Algorithm 1, we can mitigate the risk of privacy leakage. Figure 4 shows that MaCoDE allows for a trade-off between synthetic data quality (feature selection performance) and privacy preservability (DCR, Distance to Closest Record) as $\tau$ increases (see the Appendix for detailed results).
Figure 2: Trade-off between quality and privacy. Left: feature selection performance. Right: DCR. Error bars represent standard errors. See the Appendix for detailed results.
Theoretical Results
In this section, we aim to provide theoretical insights into MaCoDE’s capabilities in conditional density estimation. Firstly, we consider a factorization of the ground-truth joint PDF $p^{*}({\bf x}_{1},\cdots,{\bf x}_{p})$ according to an arbitrary permutation $\sigma=[\sigma_{1},\sigma_{2},\cdots,\sigma_{p}]$ of the indices $\{1,2,\cdots,p\}$:
$p^{*}_{\sigma_{1}}({\bf x}_{\sigma_{1}})\cdot p^{*}_{\sigma_{2}}({\bf x}_{\sigma_{2}}|{\bf x}_{\sigma_{1}})\cdot p^{*}_{\sigma_{3}}({\bf x}_{\sigma_{3}}|{\bf x}_{\sigma_{1}},{\bf x}_{\sigma_{2}})$
××$\times\cdots\times p^{*}_{\sigma_{p}}({\bf x}_{\sigma_{p}}|{\bf x}_{\sigma_{1}},\cdots,{\bf x}_{\sigma_{p-1}}).$
As discussed in Remark 1, our objective function (5) facilitates the estimation of $p_{\sigma_{j}}^{*}({\bf x}_{\sigma_{j}}|{\bf x}_{\sigma_{1}},\cdots,{\bf x}_{\sigma_{j-1}})$ for any $\sigma$ and $j$ by utilizing the mask distribution of Definition 2.
Without loss of generality, in this section, let $\sigma$ be an identity permutation. Since $c^{*}_{j}$ is the conditional density of $\hat{F}_{j}({\bf x}_{j})$, the conditional probability of $\hat{F}_{j}({\bf x}_{j})$ in the $l$th bin, $\pi_{jl}^{*}({\bf x}_{-j})$, is defined as
[…]
where ${\bf x}_{-j}\coloneqq({\bf x}_{1},\cdots,{\bf x}_{j-1})$.
Assumption 2.
For all $j$, there exists $K_{j}\geq 0$ such that $c^{*}_{j}$ is $K_{j}$-Lipschitz.
Assumption 2 implies that $c^{*}_{j}$ is constrained in its rate of change, which allows us to estimate the conditional density using a piece-wise constant function (Wasserman 2006; Tsybakov and Tsybakov 2009).
Based on (2), by the change of variable under Assumption 1, we define our conditional density estimator for $p^{*}_{j}$ as follows:
$\hat{p}_{j}({\bf x}_{j}|{\bf x}_{-j};\theta)=\hat{c}_{j}(\hat{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)\cdot\hat{p}({\bf x}_{j})$ (6)
$=$ $\left(\sum_{l=1}^{L}\frac{\mathbb{I}(\hat{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))}{1/L}\cdot\pi_{jl}\right)\cdot\hat{p}({\bf x}_{j}),$
where $\pi_{jl}$ denotes $\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta)$ and ${\bf m}^{(j)}$ is a masking vector such that
${\bf m}^{(j)}_{k}$$\coloneqq$ $\begin{cases}1,&\text{if $k\in\{1,\cdots,j-1\}$}\\ 0,&\text{otherwise}\end{cases}.$
Note that $\hat{c}_{j}$ is the histogram-based conditional density estimator. Synthetic samples can be generated from (6) as follows: Firstly, perform histogram sampling from $\hat{c}_{j}$ by choosing a histogram bin according to $\pi_{jl}$ and uniformly sampling from that bin interval. Then, apply inverse transform sampling of $\hat{p}_{j}$ with respect to the output of histogram sampling. This procedure is outlined in Algorithm 1.
Proposition 1.
Under Assumption 1 and 2,
$\text{TV}\Big{(}p^{*}_{j}(\cdot|{\bf x}_{-j}),\hat{p}_{j}(\cdot|{\bf x}_{-j};\theta)\Big{)}$$\leq$ $\frac{K_{j}}{2L}+\frac{\sqrt{Bias(\theta)}}{\sqrt{2}/L}$
for all $j\in I_{C}$, and ${\bf x}\in\mathbb{R}^{p}$. Here, $\text{TV}(\cdot,\cdot)$ denotes the total variation distance, and $Bias(\theta)$ is defined as:
$Bias(\theta)=\sum_{l=1}^{L}\pi_{jl}^{*}({\bf x}_{-j})\log\pi_{jl}^{*}({\bf x}_{-j})$
-- $\mathbb{E}_{{\bf y}_{j}|{\bf x}_{-j}}\left[\sum_{l=1}^{L}\mathbb{I}({\bf y}_{j}=l)\log\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta)\right]$
where ${\bf y}_{j}|{\bf x}_{-j}$ is a random variable having a categorical distribution such that $\Pr({\bf y}_{j}=l|{\bf x}_{-j})=\Pr(g({\bf x};\hat{F})_{j}=l|{\bf x}_{-j})=\pi_{jl}^{*}({\bf x}_{-j})$ for all $l\in[L]$.
In the definition of $Bias(\theta)$, the second term on the right-hand side corresponds to the classification loss with respect to the target distribution given by (3). This implies that $Bias(\theta)$ can be minimized when the classification loss is minimized. And Proposition 1 demonstrates that the total variation distance between the ground-truth conditional density and our conditional density estimator $\hat{p}_{j}$ is upper bounded by $Bias(\theta)$.
Therefore, minimizing the orderless multi-class classification loss leads to minimizing the total variation distance between conditional distributions, making Algorithm 1 capable of generating theoretically valid synthetic samples under certain assumptions. Note that Proposition 1 holds for any arbitrary permutation $\sigma$, and the CDF transformation is crucial for our proposed method, as it ensures the validity of Proposition 1.
With Missing Data
Suppose the pattern of missingness varies individually for each observation and is defined by a corresponding missing indicator, denoted as ${\bf r}$, where the indicators for all observations are represented as $\{{\bf r}^{(i)}\}_{i=1}^{n}$. Here, ${\bf r}_{j}=0$ indicates that ${\bf x}_{j}$ is missing, while ${\bf r}_{j}=1$ denotes that ${\bf x}_{j}$ is observed. In cases where the training data contains missing entries (e.g., not a number), it is not feasible to input these missing entries and minimize their log-likelihoods. Therefore, to handle missing value inputs using $g$, irrespective of the value of ${\bf x}_{j}$, we further define $g$ for Definition 1 as follows: for any $F$, $g({\bf x};F)_{j}\coloneqq 0$ if ${\bf r}_{j}=0$. This indicates the bin index ‘0’ is assigned to a missing value input.
Therefore, the objective function with missing data is $\min_{\theta}\sum_{i=1}^{n}\mathbb{E}_{p({\bf m})}\big{[}{\cal L}^{*}({\bf x}^{(i)},{\bf r}^{(i)},{\bf m};\theta)\big{]}$, where
${\cal L}^{*}({\bf x},{\bf r},{\bf m};\theta)$$\coloneqq$ $-\sum_{j:{\bf m}_{j}=0,{\bf r}_{j}=1}\sum_{l=1}^{L}\mathbb{I}(g({\bf x};\hat{F})_{j}=l)$
×$\times\log\pi_{jl}(g({\bf x};\hat{F})\odot\min({\bf m},{\bf r});\theta),$
$\min({\bf m},{\bf r})$ is element-wise minimum operation, and $\hat{F}$ is estimated using the observed dataset (Chenouri, Mojirsheibani, and Montazeri 2009). Note that, in ${\cal L}^{*}({\bf x},{\bf r},{\bf m};\theta)$, we do not minimize the negative log-likelihood of missing entries.
Proposition 2.
Assuming ${\bf m}\perp\!\!\!\perp{\bf r}|{\bf x}$ and ${\bf m}\perp\!\!\!\perp{\bf x}$, if the data of ${\bf x}$ is MAR, then the following holds for the missing data model $p({\bf m},{\bf r}|{\bf x})$:
$p({\bf m},{\bf r}|{\bf x})=p({\bf m},{\bf r}|{\bf x}_{obs}),$
where ${\bf x}_{obs}$ represents the observed covariates from ${\bf x}$, and the missingness pattern of $\min({\bf m},{\bf r})$ also follows the MAR mechanism.
In ${\cal L}^{*}({\bf x},{\bf r},{\bf m};\theta)$, the missingness pattern is described by $\min({\bf m},{\bf r})$, indicating that the missing data model is determined by the joint distribution $p({\bf m},{\bf r}|{\bf x})$. Thus, as the masking vector defined in Definition 2 satisfies the conditions in Proposition 2, the missingness pattern of our proposed model also conforms to the MAR mechanism according to Proposition 2 if the given data is MAR. It implies that our inference strategies based on the observed dataset can be justified (Mattei and Frellsen 2019). However, we empirically demonstrate in Section 4 that our proposed model is also applicable to other missing data scenarios.
4 Experiments
Overview
We conduct experiments in which we can provide answers to the following three experimental questions:
  • Q1.
    Does MaCoDE achieve state-of-the-art performance in synthetic data generation?
  • Q2.
    Can MaCoDE generate high-quality synthetic data even when faced with missing data scenarios?
  • Q3.
    Is MaCoDE capable of supporting multiple imputations for deriving statistically valid inferences from missing data?
Statistical fidelity Machine learning utility
Model KL $\downarrow$ GoF $\downarrow$ MMD $\downarrow$ WD $\downarrow$ SMAPE $\downarrow$ $F_{1}$uparrow Model $\uparrow$ Feature $\uparrow$
Baseline .$016_{\pm.002}.$$029_{\pm.002}.$$002_{\pm.000}.$ $1.019_{\pm.156}$$107_{\pm.008}.$$686_{\pm.023}.$$887_{\pm.018}.$$956_{\pm.005}.$
CTGAN .$221_{\pm.014}.$$561_{\pm.046}.$$094_{\pm.007}.$ $6.435_{\pm 1.011}$$256_{\pm.016}.$$411_{\pm.027}.$$208_{\pm.048}.$$417_{\pm.043}.$
TVAE .$066_{\pm.003}.$$119_{\pm.005}.$$016_{\pm.001}.$ $\underline{1.631}_{\pm.173}$$192_{\pm.011}.$$608_{\pm.021}.$$486_{\pm.041}.$$747_{\pm.027}.$
CTAB-GAN .$116_{\pm.008}.$$196_{\pm.025}.$$044_{\pm.004}.$ $3.327_{\pm.460}$$218_{\pm.012}.$$524_{\pm.026}.$$263_{\pm.042}.$$568_{\pm.041}.$
CTAB-GAN+ .$136_{\pm.018}.$$144_{\pm.010}.$$054_{\pm.007}.$ $3.971_{\pm.772}$$226_{\pm.017}.$$530_{\pm.020}.$$227_{\pm.048}.$$601_{\pm.041}.$
DistVAE .$059_{\pm.007}.$$\underline{.070}_{\pm.004}$$016_{\pm.001}.$ $2.272_{\pm.282}$$226_{\pm.017}.$$588_{\pm.021}.$$194_{\pm.048}.$$695_{\pm.030}.$
TabDDPM .$696_{\pm.117}.$$374_{\pm.087}.$$057_{\pm.011}.$ $42.916_{\pm 8.127}$$\underline{.161}_{\pm.011}$$576_{\pm.022}.$$507_{\pm.039}.$$\underline{.770}_{\pm.027}$
TabMT .$\mathbf{.011}_{\pm.001}$$\mathbf{.035}_{\pm.003}$$\underline{.012}_{\pm.001}$ $2.299_{\pm.346}$$188_{\pm.013}.$$\underline{.622}_{\pm.024}$$\underline{.528}_{\pm.039}$$761_{\pm.028}.$
MaCoDE .$\underline{.034}_{\pm.004}$$072_{\pm.004}.$$\mathbf{.007}_{\pm.001}$ $\mathbf{1.630}_{\pm.245}$$\mathbf{.158}_{\pm.010}$$\mathbf{.635}_{\pm.023}$$\mathbf{.599}_{\pm.035}$$\mathbf{.925}_{\pm.007}$
MaCoDE(MCAR) - - - - .$168_{\pm.011}.$$623_{\pm.023}.$
MaCoDE(MAR) - - - - .$167_{\pm.011}.$$626_{\pm.023}.$
MaCoDE(MNARL) - - - - .$169_{\pm.011}.$$624_{\pm.023}.$
MaCoDE(MNARQ) - - - - .$164_{\pm.010}.$$630_{\pm.023}.$
Table 1: Q1 and Q2. The means and the standard errors of the mean across 10 datasets and 10 repeated experiments are reported. Across all missingness patterns, a missingness rate of 0.3 is employed. $\uparrow$ ($\downarrow$) denotes higher (lower) is better. The best value is bolded, and the second best is underlined.
Datasets. Similar to several recent studies (Gulati and Roysdon 2023; Kotelnikov et al. 2023), we utilize 10 publicly available real tabular UCI and Kaggle [^1] https://archive.ics.uci.edu/, https://www.kaggle.com/datasets/ datasets of varying sizes and the number of columns. Detailed statistics of these datasets are provided in the Appendix. Note that we include covtype dataset, which comprises approximately 580K rows, to demonstrate the scalability of our proposed model.
Baseline models. For MaCoDE, we set $L=50$ and $\tau=1$ for all datasets. Detailed hyperparameter settings are provided in the Appendix. [^2] We run experiments using NVIDIA A10 GPU, and our experimental codes are available with pytorch. For Q1 and Q2, we compare MaCoDE with CTGAN (Xu et al. 2019), TVAE (Xu et al. 2019), CTAB-GAN (Zhao et al. 2021), CTAB-GAN+ (Zhao et al. 2023b), DistVAE (An and Jeon 2023), TabDDPM (Kotelnikov et al. 2023), and TabMT (Gulati and Roysdon 2023). For Q3, we selected the following multiple imputation models that can handle mixed-type tabular datasets: MICE (van Buuren and Groothuis-Oudshoorn 2011), GAIN (Yoon, Jordon, and van der Schaar 2018), missMDA (Josse and Husson 2016), VAEAC (Ivanov, Figurnov, and Vetrov 2019), MIWAE (Mattei and Frellsen 2019), not-MIWAE (Ipsen, Mattei, and Frellsen 2021), and EGC (Zhao, Townsend, and Udell 2022). Detailed experimental settings for these baseline models are provided in the Appendix.
To assist in interpreting the metrics, we include a baseline synthetic dataset where the synthetic data comprises half of the real training dataset. This dataset, referred to as ‘Baseline,’ serves as a soft upper bound for evaluating the quality of the synthetic data.
Additional evaluations. Due to space constraints, (1) the results for controlling data privacy levels with varying temperature parameter $\tau$, and (2) the results for Q3 (including related works and sensitivity analysis regarding the missingness rate) are provided in the Appendix.
Evaluation Metrics
For all metrics, we report the mean and standard error of the mean (error bars) across 10 different random seeds and 10 datasets. For each random seed, the dataset is randomly split into training and testing sets with an 80% training and 20% testing ratio in the evaluation of Q1 and Q2. During evaluation, the synthetic dataset is generated to have the same number of samples as the real training dataset.
Q1. To evaluate the quality of generated synthetic data, we employ two metrics: statistical fidelity (Qian, Davis, and van der Schaar 2023) and machine learning utility (Hansen et al. 2023). For statistical fidelity, we utilize the Kullback–Leibler divergence (KL) and the Goodness-of-Fit (GoF) test (continuous: the two-sample Kolmogorov-Smirnov test statistic, categorical: the Chi-Squared test statistic) to assess marginal distributional similarity. Additionally, we employ the Maximum Mean Discrepancy (MMD) and 1-Wasserstein distance (WD) to measure joint distributional similarity. These metrics measure how well the synthetic data maintains statistical fidelity to the real training dataset.
Regarding machine learning utility, we use four metrics outlined in (Hansen et al. 2023): regression performance (SMAPE, symmetric mean absolute percentage error), classification performance ($F_{1}$), model selection performance (Model), and feature selection performance (Feature). These metrics are assessed by fitting the machine learning model on the real training and synthetic datasets individually and then comparing the performance of the two models on the test dataset. For a detailed evaluation procedure of Q1, refer to the Appendix.
Q2. To assess our proposed model’s capability to generate high-quality synthetic datasets despite missing values in the training data, we evaluate the model trained on incomplete training datasets. Since other metrics require a complete training dataset for measurement, we only assess synthetic data quality using downstream regression and classification tasks on the test dataset, as these metrics can be evaluated using the test dataset alone, as in Q1. See the Appendix for a detailed evaluation procedure.
Q3.333Since some off-the-shelf packages (missMDA, gcimpute) fail to deliver useful results, we include only 5 datasets that provide meaningful results: abalone, banknote, breast, redwine, and whitewine. We assess the effectiveness of multiple imputations by employing interval inference for the population mean, which was proposed by Rubin (Rubin and Schenker 1986). We report the bias, coverage, and confidence interval length. The detailed evaluation procedure for multiple imputations and the missing value generation mechanisms (MCAR, MAR, MNARL, MNARQ) is provided in the Appendix.
Results
Q1. As shown in Table 1, MaCoDE consistently achieves the highest metric scores in both joint distributional similarities and machine learning utility while also achieving competitive performance in marginal distributional similarity. This underscores the effectiveness of the synthetic data generation method based on estimating conditional distribution in preserving the joint statistical fidelity of the original data and enhancing the utility of synthetic data for downstream machine learning tasks. Notably, MaCoDE demonstrates remarkable performance in the feature selection downstream task.
Remark 2 (How does MaCoDE achieve the remarkable performance in feature selection?).
We attribute the effectiveness of the feature selection downstream task to our emphasis on estimating ‘conditional’ distributions. In Random Forest (Breiman 2001), each node in a decision tree represents a conditional distribution of a variable conditioned on the splits made by the tree up to that node, and the feature importance is determined by the purity of nodes. Therefore, accurately estimating the conditional distribution can lead to higher performance in preserving the feature importance ranking.
Q2. The missing data mechanism within the parentheses refers to the mechanism applied to the training dataset on which MaCoDE was trained. Despite encountering missing data scenarios such as MAR, MCAR, MNARL, and MNARQ, Table 1 illustrates MaCoDE’s ability to generate high-quality synthetic data regarding machine learning utility while handling incomplete training datasets without significant performance degradation. Even in the presence of missing entries, MaCoDE achieves better metric scores than most baseline models in terms of SMAPE, except for TabDDPM. Additionally, concerning the $F_{1}$ score, MaCoDE either competes competitively or achieves a higher score than other baseline models.
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE .$010_{\pm.001}.$$845_{\pm.019}.$$\mathbf{.040}_{\pm.002}$
GAIN .$019_{\pm.002}.$$633_{\pm.033}.$$\mathbf{.040}_{\pm.002}$
missMDA .$015_{\pm.001}.$$700_{\pm.022}.$$043_{\pm.002}.$
VAEAC .$\underline{.008}_{\pm.001}$$905_{\pm.016}.$$\mathbf{.040}_{\pm.002}$
MIWAE .$\mathbf{.006}_{\pm.000}$$\underline{.952}_{\pm.012}$$043_{\pm.002}.$
not-MIWAE .$\mathbf{.006}_{\pm.000}$$\mathbf{.949}_{\pm.012}$$\underline{.042}_{\pm.002}$
EGC .$\mathbf{.006}_{\pm.000}$${.996}_{\pm.004}.$$058_{\pm.002}.$
MaCoDE(MAR) .$\mathbf{.006}_{\pm.000}$$963_{\pm.009}.$$051_{\pm.003}.$
Table 2: Q3 under MAR at 0.3 missingness. The means and standard errors of the mean across 5 datasets and 10 repeated experiments are reported. $\downarrow$ denotes lower is better. Coverage close to 0.95 indicates better performance. The best value is bolded, and the second best is underlined.
Q3. The missing data mechanism within the parentheses refers to the mechanism applied to the dataset on which MaCoDE was trained. Table 5 indicates that MaCoDE consistently exhibits competitive performance against all baseline models across metrics assessing multiple imputation performances, including bias, coverage, and confidence interval length. This suggests our proposed approach can support multiple imputations for deriving statistically valid inferences from missing data with the MAR mechanism.
Figure 3: Sensitivity analysis with respect to missingness rate using kings dataset is performed for Q2 under MAR. The analysis focuses on machine learning utility. Results are reported as means and standard errors of the mean from 10 repeated experiments, with error bars representing the standard errors.
Sensitivity analysis. We also conducted a sensitivity analysis by varying the missingness rate of kings dataset. Figure 3 illustrates that, in terms of SMAPE, MaCoDE maintains competitive performance even as the missingness rate increases. Concerning the $F_{1}$ score, MaCoDE outperforms other models at missingness rates of 0.1 and 0.3, but its performance declines beyond a missingness rate of 0.5.
5 Conclusions and Limitations
This paper introduces an approach to generating synthetic data for mixed-type tabular datasets. Our proposed method integrates histogram-based non-parametric conditional density estimation and the MLM-based approach while bridging the theoretical gap between distributional learning and the consecutive multi-class classification task of MLM. Although our primary goal is to generate synthetic data with high MLu, we empirically demonstrate that we achieve high joint statistical fidelity and MLu simultaneously. Furthermore, empirical experiments validate that our proposed model can generate high-quality synthetic tabular datasets in terms of MLu even when incomplete training datasets are given.
Although MaCoDE demonstrates the ability to perform ‘arbitrary’ conditional density estimation by accommodating various combinations of conditioning sets and target variables, and despite empirical results showing its effectiveness in handling diverse distributions of continuous columns and generating high-quality synthetic data, the model has theoretical limitations. Specifically, it is valid under Lipschitz continuity (Assumption 2). Addressing the limitation of accommodating a broader range of continuous distributions is an important direction for future work.
Appendix A Appendix / supplemental material
Proof of Proposition 1
Proof.
Without loss of generality, let $\sigma$ be an identity permutation. We equally space grid points on the interval $[0,1]$ such that bins are represented as $[b_{l-1},b_{l}),l=1,\cdots,L$, where $h$ denotes the bin width such that $Lh=1$. We consider arbitrary $j\in I_{C}$ and ${\bf x}\in\mathbb{R}^{p}$, where ${\bf x}$ is given and fixed. For notational simplicity, let ${\bf x}_{-j}\coloneqq({\bf x}_{1},\cdots,{\bf x}_{j-1})$.
Suppose that $\hat{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l})$. By the mean value theorem, there exists $u^{*}\in(b_{l-1},b_{l})$ such that
[…]
And we have
$\Big{|}c^{*}_{j}(\hat{F}({\bf x}_{j})|{\bf x}_{-j})-\hat{c}_{j}(\hat{F}({\bf x}_{j})|{\bf x}_{-j};\theta)\Big{|}$
$=$ $\left|c^{*}_{j}(\hat{F}({\bf x}_{j})|{\bf x}_{-j})-\frac{\pi_{jl}(g({\bf x};F)\odot{\bf m}^{(j)};\theta)}{h}\right|$
$\leq$ $\left|c^{*}_{j}(\hat{F}({\bf x}_{j})|{\bf x}_{-j})-c^{*}_{j}(u^{*}|{\bf x}_{-j})\right|+\left|c^{*}_{j}(u^{*}|{\bf x}_{-j})-\frac{\pi_{jl}(g({\bf x};F)\odot{\bf m}^{(j)};\theta)}{h}\right|$
$\leq$ $|\hat{F}({\bf x}_{j})-u^{*}|\left|\frac{c^{*}_{j}(\hat{F}({\bf x}_{j})|{\bf x}_{-j})-c^{*}_{j}(u^{*}|{\bf x}_{-j})}{\hat{F}({\bf x}_{j})-u^{*}}\right|+\left|\frac{\pi_{jl}^{*}({\bf x}_{-j})}{h}-\frac{\pi_{jl}(g({\bf x};F)\odot{\bf m}^{(j)};\theta)}{h}\right|$
[…]
$\leq$ $hK_{j}+\frac{1}{h}\sum_{l=1}^{L}\left|\pi_{jl}^{*}({\bf x}_{-j})-\pi_{jl}(g({\bf x};F)\odot{\bf m}^{(j)};\theta)\right|.$
where the first equality follows from the definition of $\hat{c}_{j}$.
By the Pinsker’s inequality,
$\frac{1}{2}\sum_{l=1}^{L}\left|\pi_{jl}^{*}({\bf x}_{-j})-\pi_{jl}(g({\bf x};F)\odot{\bf m}^{(j)};\theta)\right|$
$\leq$ $\frac{1}{\sqrt{2}}\left(\sum_{l=1}^{L}\pi_{jl}^{*}({\bf x}_{-j})\log\pi_{jl}^{*}({\bf x}_{-j})-\sum_{l=1}^{L}\pi_{jl}^{*}({\bf x}_{-j})\log\pi_{jl}(g({\bf x};F)\odot{\bf m}^{(j)};\theta)\right)^{1/2}$
$\leq$ $\frac{1}{\sqrt{2}}\left(\sum_{l=1}^{L}\pi_{jl}^{*}({\bf x}_{-j})\log\pi_{jl}^{*}({\bf x}_{-j})-\mathbb{E}_{{\bf y}_{j}|{\bf x}_{-j}}\left[\sum_{l=1}^{L}\mathbb{I}({\bf y}_{j}=l)\log\pi_{jl}(g({\bf x};F)\odot{\bf m}^{(j)};\theta)\right]\right)^{1/2},$
where ${\bf y}_{j}|{\bf x}_{-j}$ is a random variable having a categorical distribution such that $\Pr({\bf y}_{j}=l|{\bf x}_{-j})=\Pr(g({\bf x};\hat{F})_{j}=l|{\bf x}_{-j})=\pi_{jl}^{*}({\bf x}_{-j})$ for all $l\in[L]$.
Then,
$\Big{|}c^{*}_{j}(\hat{F}({\bf x}_{j})|{\bf x}_{-j})-\hat{c}_{j}(\hat{F}({\bf x}_{j})|{\bf x}_{-j};\theta)\Big{|}$ (7)
$\leq$ $hK_{j}+\frac{\sqrt{2}}{h}\left(\sum_{l=1}^{L}\pi_{jl}^{*}({\bf x}_{-j})\log\pi_{jl}^{*}({\bf x}_{-j})-\mathbb{E}_{{\bf y}_{j}|{\bf x}_{-j}}\left[\sum_{l=1}^{L}\mathbb{I}({\bf y}_{j}=l)\log\pi_{jl}(g({\bf x};F)\odot{\bf m}^{(j)};\theta)\right]\right)^{1/2}.$
The total variation distance between $p^{*}_{j}(\cdot|{\bf x}_{-j})$ and $\hat{p}_{j}(\cdot|{\bf x}_{-j};\theta)$ is written as
$\text{TV}\Big{(}p^{*}_{j}(\cdot|{\bf x}_{-j}),\hat{p}_{j}(\cdot|{\bf x}_{-j};\theta)\Big{)}$
$=$ $\frac{1}{2}\int_{\mathbb{R}}\Big{|}p^{*}_{j}({\bf x}_{j}|{\bf x}_{-j})-\hat{p}_{j}({\bf x}_{j}|{\bf x}_{-j};\theta)\Big{|}d{\bf x}_{j}$
[…]
$\leq$ $\frac{hK_{j}}{2}+\frac{1}{\sqrt{2}h}\left(\sum_{l=1}^{L}\pi_{jl}^{*}({\bf x}_{-j})\log\pi_{jl}^{*}({\bf x}_{-j})-\mathbb{E}_{{\bf y}_{j}|{\bf x}_{-j}}\left[\sum_{l=1}^{L}\mathbb{I}({\bf y}_{j}=l)\log\pi_{jl}(g({\bf x};F)\odot{\bf m}^{(j)};\theta)\right]\right)^{1/2}$
$=$ $\frac{hK_{j}}{2}+\frac{\sqrt{Bias(\theta)}}{\sqrt{2}h}$
$=$ $\frac{K_{j}}{2L}+\frac{\sqrt{Bias(\theta)}}{\sqrt{2}/L},$
where the inequality holds for (7).
The proof is complete. ∎
Proof of Proposition 2
Proof.
Since the data is MAR, $p({\bf r}|{\bf x})=p({\bf r}|{\bf x}_{obs})$ and
$p({\bf m},{\bf r}|{\bf x})$ $=$ $p({\bf r}|{\bf x})\cdot p({\bf m}|{\bf x})=p({\bf r}|{\bf x})\cdot p({\bf m})$
$=$ $p({\bf r}|{\bf x}_{obs})\cdot p({\bf m})=p({\bf m},{\bf r}|{\bf x}_{obs}).$
The proof is complete. ∎
Dataset Descriptions
Download links.
  • abalone (Nash et al. 1995): https://archive.ics.uci.edu/dataset/1/abalone
  • banknote (Lohweg 2013): https://archive.ics.uci.edu/dataset/267/banknote+authentication
  • breast (Wolberg et al. 1995):
    https://archive.ics.uci.edu/dataset/17/breast+cancer+wisconsin+diagnostic
  • concrete (Yeh 2007):
    https://archive.ics.uci.edu/dataset/165/concrete+compressive+strength
  • covertype (Blackard 1998): https://www.kaggle.com/datasets/uciml/forest-cover-type-dataset
  • kings (CC0: Public Domain): https://www.kaggle.com/datasets/harlfoxem/housesalesprediction
  • letter (Slate 1991): https://archive.ics.uci.edu/dataset/59/letter+recognition
  • loan (CC0: Public Domain): https://www.kaggle.com/datasets/teertha/personal-loan-modeling
  • redwine (Cortez et al. 2009): https://archive.ics.uci.edu/dataset/186/wine+quality
  • whitewine (Cortez et al. 2009): https://archive.ics.uci.edu/dataset/186/wine+quality
Dataset Train/Test Split #continuous #categorical Classification Target
abalone 3.3K/0.8K 7 2 Rings
banknote 1.1K/0.3K 4 1 class
breast 0.5K/0.1K 30 1 Diagnosis
concrete 0.8K/0.2K 8 1 Age
covtype 0.46M/0.12M 10 1 Cover_Type
kings 17.3K/4.3K 11 7 grade
letter 16K/4K 16 1 lettr
loan 4K/1K 5 6 Personal Loan
redwine 1.3K/0.3K 11 1 quality
whitewine 3.9K/1K 11 1 quality
Table 3: Description of datasets. #continuous represents the number of continuous and ordinal variables. #categorical denotes the number of categorical variables. The ‘Classification Target’ refers to the variable used as the response variable in a classification task to evaluate machine learning utility.
Experimental Settings for Reproduction
  • We run experiments using NVIDIA A10 GPU, and our experimental codes are available with pytorch.
  • In practice, for all $j\in I_{C}$, we estimate $\hat{F}_{j}$ using empirical measure. In the presence of missing data, $\hat{F}_{j}$ is estimated solely using the observed dataset (refer to (Chenouri, Mojirsheibani, and Montazeri 2009)).
Hyper-parameters of MaCoDE: As shown in Section A, MaCoDE consistently generates high-quality synthetic data without requiring an extensive hyperparameter tuning process, unlike methods such as (Kotelnikov et al. 2023; Lee, Kim, and Park 2023; Kim, Lee, and Park 2023). This demonstrates the generalizability of our proposed model to various tabular datasets. Our implementation codes for the proposed model, MaCoDE, are provided in the supplementary material. For all tabular datasets, we applied the following hyperparameters uniformly without any additional tuning:
  • epochs: 500
  • batch size: 1024
  • learning rate: 0.001 (with AdamW optimizer (Loshchilov and Hutter 2017) with 0.001 weight decay parameter)
  • the number of bins: $L=50$
  • Transformer encoder dimension: 128
  • Transformer encoder #heads: 4
  • Transformer encoder #layer: 2
  • Transformer encoder dropout ratio: 0.0
Details of Implementing Baseline Models
To compare our proposed method with baseline models, we performed experiments in a manner similar to (Zhang et al. 2024). We adjusted the hidden or latent dimensions across various methods, ensuring that the number of trainable parameters is comparable. Since each model employs distinct neural network architectures, it would be misleading to compare their performances using the same network configuration. It’s worth noting that model size typically reflects the performance of the baseline models. Under these conditions, we reproduced the baseline methods using their official codes and default configurations (except for the hidden or latent dimensions). Our reproduced codes for the baseline models are provided in the supplementary material. Below are the detailed implementations of the baseline methods:
  • CTGAN and TVAE (Xu et al. 2019): We follow the implementations provided in the official repository [^4] https://github.com/sdv-dev/CTGAN for CTGAN and TVAE. We adopt the default hyperparameters specified in the module. To ensure fairness in comparison by aligning model sizes, we adjust the latent dimension of CTGAN and TVAE to 100.
  • CTAB-GAN and CTAB-GAN+ (Zhao et al. 2021, 2023b): We follow the implementations provided in the official repository [^5] https://github.com/Team-TUD/CTAB-GAN, https://github.com/Team-TUD/CTAB-GAN-Plus. The latent dimension is set to 100, and the maximum number of clusters is set to 10. For the specification of feature types, we use three categories: continuous, integer, and categorical variables.
  • DistVAE (An and Jeon 2023): We follow the implementations provided in the official repository [^6] https://github.com/an-seunghwan/DistVAE. We changed the latent dimension to 100 to increase the model size.
  • TabDDPM (Kotelnikov et al. 2023): We utilized the TabDDPM module in synthcity.Plugins  [^7] https://github.com/vanderschaarlab/synthcity for synthetic data generation. This module follows the implementations provided in the official repository.
  • TabMT (Gulati and Roysdon 2023): Rather than using K-means clustering, we utilize the Gaussian Mixture Model (GMM) to discretize continuous columns to preserve the original continuous domain. We determine the optimal number of clusters for the GMM, ranging from 2 to 10, based on the Bayesian Information Criterion (BIC). During the generation of continuous columns, we initially predict the component label and then sample from the selected Gaussian component.
  • MICE (van Buuren and Groothuis-Oudshoorn 2011): We employed the IterativeImputer package from Scikit-learn for multiple imputation using chained equations. Following the authors’ experiments, a max_iter range of 10 to 20 was considered sufficient for convergence, and we adopted this setting for our experiments. Additionally, to introduce randomness, we set imputation_order to random, which randomly selects a variable for imputation in each iteration. The remaining parameters were left at their default values to maintain the integrity of the MICE implementation.
  • GAIN (Yoon, Jordon, and van der Schaar 2018): We follow the implementations provided in the official repository [^8] https://github.com/jsyoon0823/GAIN/tree/master. As the paper does not explicitly discuss the separate handling of categorical and continuous variables, the code treats them simultaneously. Consequently, a rounding process is employed to handle categorical variables afterward.
  • missMDA (Josse and Husson 2016): We utilized the missMDA package [^9] https://cran.r-project.org/web/packages/missMDA/index.html in R for multiple imputation.
  • VAEAC (Ivanov, Figurnov, and Vetrov 2019): We follow the implementations provided in the official repository [^10] https://github.com/tigvarts/vaeac. The authors provided hyperparameters that adequately address both continuous and categorical variables, so we used these without further modification during model fitting.
  • MIWAE (Mattei and Frellsen 2019): The implemented MIWAE code in the official repository [^11] https://github.com/pamattei/miwae was designed for continuous variables only. To accommodate heterogeneous tabular datasets, we treated the conditional distribution of categorical columns as categorical distributions and employed cross-entropy loss for reconstruction. For comparison with not-MIWAE, we set the latent dimension to $p-1$.
  • not-MIWAE (Ipsen, Mattei, and Frellsen 2021): The implemented not-MIWAE code in the official repository [^12] https://github.com/nbip/notMIWAE also focused on continuous variables exclusively. To handle categorical variables, we made the same modifications as in MIWAE. For comparison with MIWAE, we set the latent dimension to $p-1$. Training was conducted for 100K steps, consistent with the official implementations.
  • EGC (Zhao, Townsend, and Udell 2022): We utilized the gcimpute package [^13] https://github.com/udellgroup/gcimpute/tree/master to implement EGC. The model is free of hyperparameters.
Tasks Model Description
Regression Random Forest Package: sklearn.ensemble.RandomForestRegressor, setting: random_state=0, defaulted values
Classification Logistic Regression Package: sklearn.linear_model.LogisticRegression, setting: random_state=0, max_iter=1000, defaulted values
Gaussian Navie Bayes Package: sklearn.naive_bayes.GaussianNB, setting: defaulted values
K-Nearest Neighbors Package: sklearn.neighbors.KNeighborsClassifier, setting: defaulted values
Decision Tree Package: sklearn.tree.DecisionTreeClassifier, setting: random_state=0, defaulted values
Random Forest Package: sklearn.ensemble.RandomForestClassifier, setting: random_state=0, defaulted values
Table 4: Regressor and classifier used to evaluate synthetic data quality in machine learning utility. The names of all parameters used in the description are consistent with those defined in corresponding packages.
Evaluation Procedure of Q1
Regression performance (SMAPE).
  1. 1.
    Train a synthesizer using the real training dataset.
  2. 2.
    Generate a synthetic dataset with the same size as the real training dataset.
  3. 3.
    Train a machine learning model (Random Forest regressor) using the synthetic dataset, where each continuous column serves as the regression target variable.
  4. 4.
    Assess regression prediction performance by averaging the SMAPE values from the test dataset for each Random Forest regressor trained on the continuous columns.
Classification performance ($F_{1}$).
  1. 1.
    Train a synthesizer using the real training dataset.
  2. 2.
    Generate a synthetic dataset with the same size as the real training dataset.
  3. 3.
    Train machine learning models (Logistic Regression, Gaussian Naive Bayes, K-Nearest Neighbors classifier, Decision Tree classifier, and Random Forest classifier) using the synthetic dataset (see Table 3 for the classification target variable and refer to Table 4 for detailed configuration).
  4. 4.
    Assess classification prediction performance by averaging the $F_{1}$ values from the test dataset from five different classifiers.
Model selection performance (Model).
  1. 1.
    Train a synthesizer using the real training dataset.
  2. 2.
    Generate a synthetic dataset with the same size as the real training dataset.
  3. 3.
    Train machine learning models (Logistic Regression, Gaussian Naive Bayes, K-Nearest Neighbors classifier, Decision Tree classifier, and Random Forest classifier) using both the real training dataset and the synthetic dataset (see Table 3 for the classification target variable and refer to Table 4 for detailed configuration).
  4. 4.
    Evaluate the classification performance (AUROC) of all trained classifiers on the test dataset.
  5. 5.
    Assess model selection performance by comparing the AUROC rank orderings of classifiers trained on the real training dataset and those trained on the synthetic dataset using Spearman’s Rank Correlation.
Feature selection performance (Feature).
  1. 1.
    Train a synthesizer using the real training dataset.
  2. 2.
    Generate a synthetic dataset with the same size as the real training dataset.
  3. 3.
    Train a Random Forest classifier using both the real training dataset and the synthetic dataset (see Table 3 for the classification target variable and refer to Table 4 for detailed configuration).
  4. 4.
    Determine the rank-ordering of important features for both classifiers.
  5. 5.
    Assess feature selection performance by comparing the feature importance rank orderings of classifiers trained on the real training dataset and those trained on the synthetic dataset using Spearman’s Rank Correlation.
Evaluation Procedure of Q2
Regression performance (SMAPE).
  1. 1.
    For each random seed, we generate the mask and train MaCoDE using a masked training dataset (i.e., incomplete dataset).
  2. 2.
    Generate a synthetic dataset with the same size as the real training dataset.
  3. 3.
    Train a machine learning model (Random Forest regressor) using the synthetic dataset, where each continuous column serves as the regression target variable.
  4. 4.
    Assess regression prediction performance by averaging the SMAPE values from the test dataset for each Random Forest regressor trained on the continuous columns.
Classification performance ($F_{1}$).
  1. 1.
    For each random seed, we generate the mask and train MaCoDE using a masked training dataset (i.e., incomplete dataset).
  2. 2.
    Generate a synthetic dataset with the same size as the real training dataset.
  3. 3.
    Train machine learning models (Logistic Regression, Gaussian Naive Bayes, K-Nearest Neighbors classifier, Decision Tree classifier, and Random Forest classifier) using the synthetic dataset (see Table 3 for the classification target variable and refer to Table 4 for detailed configuration).
  4. 4.
    Assess classification prediction performance by averaging the $F_{1}$ values from the test dataset from five different classifiers.
Evaluation of Q3
Related Works
Data missingness is a common challenge in research and practical analysis, categorized into three primary missing mechanisms: (1) Missing Completely at Random (MCAR), (2) Missing at Random (MAR), and (3) Missing Not at Random (MNAR).
Under the MCAR mechanism, the reason for missingness has no relationship with any data, neither observed nor unobserved. In other words, the likelihood of data being missing is equal across all observations. The primary advantage of MCAR is that it does not introduce bias into the data analysis. However, despite this advantage, data missingness can still reduce the statistical power of the study because of the reduced sample size.
MAR occurs when the probability of missingness is related to the observed data but not the unobserved missing data. Essentially, even though the data is missing, the mechanism assumes that the missingness is explainable by other variables in the dataset. That is, the missingness can be modeled and imputed using the information available in the data, allowing for more accurate analyses despite the missingness.
Lastly, if the missingness is not specified by either MCAR or MAR, it becomes MNAR. The MNAR is the most challenging mechanism, as it implies that the missingness is related to the unobserved data itself. In this case, the missing data is systematically different from the observed data, which introduces bias if not properly accounted for. For example, patients with severe symptoms may be less likely to report their health status, making their data missing. MNAR requires sophisticated statistical methods to address, as ignoring or improperly handling it can lead to biased and unreliable results.
Recent methods employ the deep generative model to estimate a joint distribution and generate samples for imputations. GAN-based imputers such as GAIN (Yoon, Jordon, and van der Schaar 2018) and MisGAN (Li, Jiang, and Marlin 2019) adopt an adversarial learning approach to generate both missing entries and masking vectors. Other approaches like VAEAC (Ivanov, Figurnov, and Vetrov 2019), HI-VAE (Nazábal et al. 2020), and ReMasker (Du, Melis, and Wang 2024) employ strategies that allow them to learn conditional distributions on arbitrary conditioning sets using the uniform masking strategy. MIWAE (Mattei and Frellsen 2019), based on the Importance Weighted Autoencoder (Burda, Grosse, and Salakhutdinov 2015), demonstrates that under mild conditions and the MAR assumption, the target likelihood can be approximated regardless of the imputation function. Additionally, not-MIWAE (Ipsen, Mattei, and Frellsen 2021) extends MIWAE to handle cases where the data is under MNAR assumption by modeling missing entries as a latent variable. In parallel, the Optimal Transport (OT)-based method employs distributional matching, utilizing the 2-Wasserstein distance to compare distributions in both data and latent spaces (Muzellec et al. 2020; Zhao et al. 2023a). To handle mixed-type tabular datasets, (Zhao, Townsend, and Udell 2022) introduced EGC (extended Gaussian copula), which relies on a latent Gaussian distribution to support single and multiple imputations.
Evaluation Metrics
Algorithm 2 Evaluation procedure for multiple imputation (van Buuren 2012; Rubin and Schenker 1986)
Input: Complete dataset $D=\{x_{i}\}_{i=1}^{n}$
Output: Bias, Coverage, and Confidence interval length
1:  Target estimand: $Q^{*}=1/n\sum_{i=1}^{n}\mathbb{I}(x_{i}>\bar{x})$, where $\bar{x}=1/n\sum_{i=1}^{n}x_{i}$
2:  for (random seeds) $s=1,\cdots,S$ do
3:     Generate missing dataset $D^{(s)}$
4:     Training the imputation model using the dataset $D^{(s)}$
5:     Perform multiple imputation: $\hat{D}_{m}^{(s)},m=1,\cdots,M$
6:     for $m=1,2,\cdots,M$ do
7:        $\hat{Q}_{m}^{(s)}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(x_{i}>\bar{x})$, where $\hat{D}_{m}^{(s)}=\{x_{i}\}_{i=1}^{n}$ and $\bar{x}=\frac{1}{n}\sum_{i=1}^{n}x_{i}$
8:        $\hat{U}_{m}^{(s)}=\hat{Q}_{m}^{(s)}(1-\hat{Q}_{m}^{(s)})/n$
9:     end for
10:     $\hat{Q}^{(s)}=\frac{1}{M}\sum_{m=1}^{M}\hat{Q}_{m}^{(s)}$
11:     $\hat{U}^{(s)}=\frac{1}{M}\sum_{m=1}^{M}\hat{U}_{m}^{(s)}+(M+1)/M\cdot\frac{1}{M-1}\sum_{m=1}^{M}(\hat{U}_{m}^{(s)}-\bar{U}^{(s)})^{2}$, where $\bar{U}^{(s)}=\frac{1}{M}\sum_{m=1}^{M}\hat{U}_{m}^{(s)}$
12:  end for
13:  Bias: $\frac{1}{S}\sum_{s=1}^{S}|\hat{Q}^{(s)}-Q^{*}|$
14:  Coverage: $\frac{1}{S}\sum_{s=1}^{S}\mathbb{I}(Q^{*}\in(\hat{Q}^{(s)}\pm 1.96\cdot\sqrt{\hat{U}^{(s)}}))$
15:  Confidence interval length: $\frac{1}{S}\sum_{s=1}^{S}2\cdot 1.96\cdot\sqrt{\hat{U}^{(s)}}$
For Q3, we selected the following multiple imputation models that can handle heterogeneous tabular datasets: MICE (van Buuren and Groothuis-Oudshoorn 2011), GAIN (Yoon, Jordon, and van der Schaar 2018), missMDA (Josse and Husson 2016), VAEAC (Ivanov, Figurnov, and Vetrov 2019), MIWAE (Mattei and Frellsen 2019), not-MIWAE (Ipsen, Mattei, and Frellsen 2021), and EGC (Zhao, Townsend, and Udell 2022).
We assess the effectiveness of multiple imputations by employing interval inference for the population mean, which was proposed by Rubin (Rubin and Schenker 1986). However, since obtaining the population mean is not feasible, we utilize the sample column-wise mean of the complete dataset as a parameter of interest (Lee and Huber 2021; Zhang et al. 2023). The evaluation procedure for multiple imputations is outlined in Algorithm 2, and we utilize torch.random.manual_seed() to set seeds in performing multiple imputations. We report the mean and standard error of bias, coverage, and confidence interval length across 10 different random seeds, all continuous columns, and datasets (van Buuren 2012). Note that we do not split the data into training and test sets (Ipsen, Mattei, and Frellsen 2021).
Remark 3.
We acknowledge that existing imputation methods have been evaluated using RMSE (root mean square error), a metric for assessing single imputation methods (Ipsen, Mattei, and Frellsen 2021; Mattei and Frellsen 2019; Nazábal et al. 2020; Muzellec et al. 2020; Zhao et al. 2023a; Jarrett et al. 2022). However, since our objective focuses on distributional learning rather than recovering missing values, we adopt the evaluation procedure proposed in (Rubin and Schenker 1986; van Buuren 2012), better suited for assessing multiple imputation methods.
Following (Muzellec et al. 2020; Jarrett et al. 2022; Zhao et al. 2023a), we generate the missing value mask for each dataset with three mechanisms in four settings. (MCAR) In the MCAR setting, each value is masked according to the realization of a Bernoulli random variable with a fixed parameter. (MAR) In the MAR setting, for each experiment, a fixed subset of variables that cannot have missing values is sampled. Then, the remaining variables have missing values according to a logistic model with random weights, which takes the non-missing variables as inputs. A bias term is fitted using line search to attain the desired proportion of missing values. (MNAR) Finally, two different mechanisms are implemented in the MNAR setting. The first, MNARL, is identical to the previously described MAR mechanism, but the inputs of the logistic model are then masked by an MCAR mechanism. Hence, the logistic model’s outcome depends on missing values. The second mechanism, MNARQ, samples a subset of variables whose values in the lower and upper $p$th percentiles are masked according to a Bernoulli random variable, and the values in-between are left not missing.
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE .$010_{\pm.001}.$$845_{\pm.019}.$$\mathbf{.040}_{\pm.002}$
GAIN .$019_{\pm.002}.$$633_{\pm.033}.$$\mathbf{.040}_{\pm.002}$
missMDA .$015_{\pm.001}.$$700_{\pm.022}.$$043_{\pm.002}.$
VAEAC .$\underline{.008}_{\pm.001}$$905_{\pm.016}.$$\mathbf{.040}_{\pm.002}$
MIWAE .$\mathbf{.006}_{\pm.000}$$\underline{.952}_{\pm.012}$$043_{\pm.002}.$
not-MIWAE .$\mathbf{.006}_{\pm.000}$$\mathbf{.949}_{\pm.012}$$\underline{.042}_{\pm.002}$
EGC .$\mathbf{.006}_{\pm.000}$${.996}_{\pm.004}.$$058_{\pm.002}.$
MaCoDE(MAR) .$\mathbf{.006}_{\pm.000}$$963_{\pm.009}.$$051_{\pm.003}.$
Table 5: Q3: Multiple imputation under MAR at 0.3 missingness. The means and standard errors of the mean across 5 datasets and 10 repeated experiments are reported. $\downarrow$ denotes lower is better. Coverage close to 0.95 indicates better performance. The best value is bolded, and the second best is underlined.
Results
The missing data mechanism within the parentheses refers to the mechanism applied to the dataset on which MaCoDE was trained. Table 5 indicates that MaCoDE consistently exhibits competitive performance against all baseline models across metrics assessing multiple imputation performances, including bias, coverage, and confidence interval length. This suggests our proposed approach can support multiple imputations for deriving statistically valid inferences from missing data with the MAR mechanism.
Additional Experiments
Privacy Preservability
  • Evaluation metrics: The $k$-anonymity property (Sweeney 2002) is a measure used to assess the level of privacy protection in synthetic data. It ensures that each individual’s information in the dataset cannot be distinguished from that of at least $k-1$ other individuals. In other words, each record in the dataset has at least $k-1$ similar records in terms of quasi-identifiers, attributes that could potentially identify a subject. A higher $k$ value implies a higher level of anonymity and better privacy preservation.
    DCR (Distance to Closest Record) (Park et al. 2018; Zhao et al. 2021) is defined as the distances between all real training and synthetic samples. A higher DCR value indicates more effective privacy preservation, indicating a lack of overlap between the real training data and the synthetic samples. Conversely, an excessively large DCR score suggests a lower quality of the generated synthetic dataset. Therefore, the DCR metric provides insights into both the privacy-preserving capability and the quality of the synthetic dataset.
    Attribute disclosure (Choi et al. 2017; Matwin et al. 2015) refers to the situation where attackers can uncover additional covariates of a record by leveraging a subset of covariates they already possess, along with similar records from the synthetic dataset. To quantify the extent to which attackers can accurately identify these additional covariates, we employ classification metrics. Higher attribute disclosure metrics indicate an increased risk of privacy leakage, implying that attackers can precisely infer unknown variables. In terms of privacy concerns, attribute disclosure can be considered a more significant issue than membership inference attacks, as attackers are assumed to have access to only a subset of covariates for a given record.
  • Evaluation procedure: We evaluate the $k$-anonymity following the approach described in (Qian, Davis, and van der Schaar 2023)  [^14] https://github.com/vanderschaarlab/synthcity. Additionally, similar to (Zhao et al. 2021), we define DCR as the $5^{th}$ percentile of the $L_{2}$ distances between all real training samples and synthetic samples. Since DCR relies on $L_{2}$ distance and continuous variables, it is computed solely using continuous variables. We assess attribute disclosure using the methodology outlined in (Choi et al. 2017).
  • Result: The right panel of Figure 4 demonstrates MaCoDE’s capability to regulate the privacy level by adjusting the temperature parameter $\tau$. Simultaneously, the left panel illustrates that the quality of synthetic data, measured by feature selection performance, remains notable even with increasing privacy levels from $\tau=1$ to $\tau=3$. However, increasing $\tau$ beyond 3 leads to declining feature selection performance compared to other models despite DCR remaining competitive. For additional results on other metrics related to the trade-off between privacy level and synthetic data quality as $\tau$ varies, please refer to Table 6 and Table 7.
Figure 4: Trade-off between privacy and quality. Left: feature selection performance (synthetic data quality). Right: DCR (privacy preservability). The means and standard errors of the mean across 10 datasets and 10 repeated experiments are reported. Error bars represent the standard errors of the mean.
Statistical fidelity Machine learning utility
Model KL $\downarrow$ GoF $\downarrow$ MMD $\downarrow$ WD $\downarrow$ SMAPE $\downarrow$ $F_{1}$uparrow Model $\uparrow$ Feature $\uparrow$
Baseline .$016_{\pm.002}.$$029_{\pm.002}.$$002_{\pm.000}.$ $1.019_{\pm.156}$$107_{\pm.008}.$$686_{\pm.023}.$$887_{\pm.018}.$$956_{\pm.005}.$
CTGAN .$221_{\pm.014}.$$561_{\pm.046}.$$094_{\pm.007}.$ $6.435_{\pm 1.011}$$256_{\pm.016}.$$411_{\pm.027}.$$208_{\pm.048}.$$417_{\pm.043}.$
TVAE .$066_{\pm.003}.$$119_{\pm.005}.$$016_{\pm.001}.$ $1.631_{\pm.173}$$192_{\pm.011}.$$608_{\pm.021}.$$486_{\pm.041}.$$747_{\pm.027}.$
CTAB-GAN .$116_{\pm.008}.$$196_{\pm.025}.$$044_{\pm.004}.$ $3.327_{\pm.460}$$218_{\pm.012}.$$524_{\pm.026}.$$263_{\pm.042}.$$568_{\pm.041}.$
CTAB-GAN+ .$136_{\pm.018}.$$144_{\pm.010}.$$054_{\pm.007}.$ $3.971_{\pm.772}$$226_{\pm.017}.$$530_{\pm.020}.$$227_{\pm.048}.$$601_{\pm.041}.$
DistVAE .$059_{\pm.007}.$$070_{\pm.004}.$$016_{\pm.001}.$ $2.272_{\pm.282}$$226_{\pm.017}.$$588_{\pm.021}.$$194_{\pm.048}.$$695_{\pm.030}.$
TabDDPM .$696_{\pm.117}.$$374_{\pm.087}.$$057_{\pm.011}.$ $42.916_{\pm 8.127}$$161_{\pm.011}.$$576_{\pm.022}.$$507_{\pm.039}.$$770_{\pm.027}.$
TabMT .$011_{\pm.001}.$$035_{\pm.003}.$$012_{\pm.001}.$ $2.299_{\pm.346}$$188_{\pm.013}.$$622_{\pm.024}.$$528_{\pm.039}.$$761_{\pm.028}.$
MaCoDE($\tau=1$) .$034_{\pm.004}.$$072_{\pm.004}.$$007_{\pm.001}.$ $1.630_{\pm.245}$$158_{\pm.010}.$$635_{\pm.023}.$$599_{\pm.035}.$$925_{\pm.007}.$
MaCoDE($\tau=2$) .$047_{\pm.004}.$$117_{\pm.009}.$$021_{\pm.002}.$ $2.450_{\pm.302}$$177_{\pm.012}.$$601_{\pm.024}.$$432_{\pm.038}.$$871_{\pm.011}.$
MaCoDE($\tau=3$) .$071_{\pm.006}.$$290_{\pm.046}.$$035_{\pm.003}.$ $3.191_{\pm.354}$$196_{\pm.013}.$$562_{\pm.024}.$$305_{\pm.041}.$$761_{\pm.020}.$
MaCoDE($\tau=4$) .$088_{\pm.008}.$$555_{\pm.109}.$$043_{\pm.003}.$ $3.601_{\pm.383}$$210_{\pm.014}.$$528_{\pm.024}.$$128_{\pm.041}.$$639_{\pm.035}.$
MaCoDE($\tau=5$) .$099_{\pm.010}.$$847_{\pm.181}.$$045_{\pm.003}.$ $3.822_{\pm.402}$$219_{\pm.014}.$$501_{\pm.025}.$$065_{\pm.047}.$$555_{\pm.044}.$
Table 6: Trade-off between privacy and quality (statistical fidelity and machine learning utility). The means and standard errors of the mean across 10 datasets and 10 repeated experiments are reported. ‘Baseline’ refers to the result obtained using half of the real training dataset. $\uparrow$ ($\downarrow$) denotes higher (lower) is better.
Model $k$-anonimity(%) $\uparrow$ DCR $\uparrow$ AD $\downarrow$
Baseline $1.151_{\pm 0.089}$ $0.501_{\pm 0.051}$ $0.631_{\pm 0.023}$
CTGAN $2.413_{\pm 0.139}$ $1.355_{\pm 0.152}$ $0.398_{\pm 0.024}$
TVAE $1.648_{\pm 0.116}$ $0.709_{\pm 0.056}$ $0.583_{\pm 0.023}$
CTAB-GAN $2.048_{\pm 0.126}$ $0.951_{\pm 0.076}$ $0.514_{\pm 0.025}$
CTAB-GAN+ $2.215_{\pm 0.121}$ $1.016_{\pm 0.113}$ $0.471_{\pm 0.018}$
DistVAE $2.345_{\pm 0.136}$ $0.932_{\pm 0.071}$ $0.540_{\pm 0.019}$
TabDDPM $1.249_{\pm 0.108}$ $1.779_{\pm 0.391}$ $0.567_{\pm 0.024}$
TabMT $1.656_{\pm 0.135}$ $0.862_{\pm 0.072}$ $0.566_{\pm 0.021}$
MaCoDE($\tau=1$) $1.405_{\pm 0.104}$ $0.658_{\pm 0.067}$ $0.589_{\pm 0.022}$
MaCoDE($\tau=2$) $1.656_{\pm 0.109}$ $0.797_{\pm 0.070}$ $0.547_{\pm 0.022}$
MaCoDE($\tau=3$) $1.959_{\pm 0.147}$ $0.919_{\pm 0.077}$ $0.508_{\pm 0.022}$
MaCoDE($\tau=4$) $2.075_{\pm 0.159}$ $1.000_{\pm 0.082}$ $0.476_{\pm 0.022}$
MaCoDE($\tau=5$) $2.262_{\pm 0.166}$ $1.043_{\pm 0.085}$ $0.457_{\pm 0.022}$
Table 7: Privacy preservability. The means and standard errors of the mean across 10 datasets and 10 repeated experiments are reported. ‘Baseline’ refers to the result obtained using half of the real training dataset. $\uparrow$ ($\downarrow$) denotes higher (lower) is better.
Dataset abalone
Model $k$-anonimity(%) $\uparrow$ DCR $\uparrow$ AD $\downarrow$
CTGAN $3.634_{\pm 0.318}$ $0.423_{\pm 0.012}$ $0.346_{\pm 0.009}$
TVAE $2.029_{\pm 0.234}$ $0.333_{\pm 0.010}$ $0.385_{\pm 0.018}$
CTAB-GAN $2.769_{\pm 0.315}$ $0.449_{\pm 0.027}$ $0.312_{\pm 0.023}$
CTAB-GAN+ $2.718_{\pm 0.285}$ $0.311_{\pm 0.010}$ $0.323_{\pm 0.010}$
DistVAE $2.074_{\pm 0.149}$ $0.316_{\pm 0.006}$ $0.349_{\pm 0.009}$
TabDDPM $0.233_{\pm 0.047}$ $0.145_{\pm 0.003}$ $0.371_{\pm 0.016}$
TabMT $1.595_{\pm 0.617}$ $0.423_{\pm 0.007}$ $0.337_{\pm 0.010}$
MaCoDE $1.182_{\pm 0.292}$ $0.148_{\pm 0.002}$ $0.368_{\pm 0.014}$
Dataset banknote
Model $k$-anonimity(%) $\uparrow$ DCR $\uparrow$ AD $\downarrow$
CTGAN $2.862_{\pm 0.146}$ $0.341_{\pm 0.020}$ $0.763_{\pm 0.033}$
TVAE $2.607_{\pm 0.305}$ $0.220_{\pm 0.007}$ $0.920_{\pm 0.021}$
CTAB-GAN $3.118_{\pm 0.202}$ $0.275_{\pm 0.008}$ $0.847_{\pm 0.025}$
CTAB-GAN+ $2.799_{\pm 0.151}$ $0.279_{\pm 0.012}$ $0.560_{\pm 0.038}$
DistVAE $3.081_{\pm 0.267}$ $0.290_{\pm 0.010}$ $0.757_{\pm 0.027}$
TabDDPM $0.985_{\pm 0.180}$ $0.168_{\pm 0.007}$ $0.997_{\pm 0.003}$
TabMT $2.662_{\pm 0.262}$ $0.257_{\pm 0.009}$ $0.910_{\pm 0.015}$
MaCoDE $2.552_{\pm 0.255}$ $0.076_{\pm 0.005}$ $0.930_{\pm 0.016}$
Dataset breast
Model $k$-anonimity(%) $\uparrow$ DCR $\uparrow$ AD $\downarrow$
CTGAN $1.582_{\pm 0.383}$ $5.536_{\pm 0.167}$ $0.446_{\pm 0.050}$
TVAE $0.813_{\pm 0.239}$ $2.148_{\pm 0.026}$ $0.900_{\pm 0.042}$
CTAB-GAN $0.857_{\pm 0.198}$ $2.807_{\pm 0.078}$ $0.850_{\pm 0.037}$
CTAB-GAN+ $1.473_{\pm 0.289}$ $4.025_{\pm 0.315}$ $0.596_{\pm 0.078}$
DistVAE $2.264_{\pm 0.390}$ $2.666_{\pm 0.060}$ $0.810_{\pm 0.036}$
TabDDPM $2.835_{\pm 0.301}$ $12.733_{\pm 0.896}$ $0.740_{\pm 0.089}$
TabMT $0.967_{\pm 0.164}$ $2.718_{\pm 0.054}$ $0.825_{\pm 0.041}$
MaCoDE $0.527_{\pm 0.140}$ $2.264_{\pm 0.058}$ $0.906_{\pm 0.032}$
Dataset concrete
Model $k$-anonimity(%) $\uparrow$ DCR $\uparrow$ AD $\downarrow$
CTGAN $3.750_{\pm 0.305}$ $1.236_{\pm 0.076}$ $0.156_{\pm 0.040}$
TVAE $2.464_{\pm 0.386}$ $0.681_{\pm 0.018}$ $0.340_{\pm 0.054}$
CTAB-GAN $3.519_{\pm 0.216}$ $0.973_{\pm 0.044}$ $0.310_{\pm 0.054}$
CTAB-GAN+ $4.066_{\pm 0.201}$ $0.972_{\pm 0.034}$ $0.310_{\pm 0.033}$
DistVAE $4.235_{\pm 0.231}$ $1.012_{\pm 0.037}$ $0.358_{\pm 0.051}$
TabDDPM $1.663_{\pm 0.253}$ $0.425_{\pm 0.011}$ $0.388_{\pm 0.044}$
TabMT $4.029_{\pm 0.228}$ $0.949_{\pm 0.038}$ $0.308_{\pm 0.053}$
MaCoDE $2.864_{\pm 0.293}$ $0.239_{\pm 0.019}$ $0.348_{\pm 0.043}$
Dataset covtype
Model $k$-anonimity(%) $\uparrow$ DCR $\uparrow$ AD $\downarrow$
CTGAN $0.215_{\pm 0.000}$ $0.451_{\pm 0.003}$ $0.490_{\pm 0.023}$
TVAE $0.215_{\pm 0.000}$ $0.415_{\pm 0.002}$ $0.639_{\pm 0.009}$
CTAB-GAN $0.215_{\pm 0.000}$ $0.418_{\pm 0.004}$ $0.579_{\pm 0.004}$
CTAB-GAN+ $0.215_{\pm 0.000}$ $0.353_{\pm 0.001}$ $0.645_{\pm 0.001}$
DistVAE $0.215_{\pm 0.000}$ $0.566_{\pm 0.002}$ $0.612_{\pm 0.002}$
TabDDPM $0.211_{\pm 0.002}$ $0.313_{\pm 0.001}$ $0.653_{\pm 0.002}$
TabMT $0.215_{\pm 0.000}$ $0.409_{\pm 0.001}$ $0.671_{\pm 0.003}$
MaCoDE $0.215_{\pm 0.000}$ $0.266_{\pm 0.000}$ $0.696_{\pm 0.001}$
Dataset kings
Model $k$-anonimity(%) $\uparrow$ DCR $\uparrow$ AD $\downarrow$
CTGAN $1.243_{\pm 0.168}$ $0.502_{\pm 0.006}$ $0.601_{\pm 0.004}$
TVAE $0.848_{\pm 0.066}$ $0.361_{\pm 0.003}$ $0.685_{\pm 0.003}$
CTAB-GAN $1.158_{\pm 0.103}$ $0.461_{\pm 0.004}$ $0.631_{\pm 0.003}$
CTAB-GAN+ $1.028_{\pm 0.126}$ $0.411_{\pm 0.006}$ $0.655_{\pm 0.003}$
DistVAE $0.699_{\pm 0.113}$ $0.488_{\pm 0.002}$ $0.658_{\pm 0.004}$
TabDDPM $0.236_{\pm 0.032}$ $0.325_{\pm 0.003}$ $0.668_{\pm 0.005}$
TabMT $0.396_{\pm 0.103}$ $0.427_{\pm 0.003}$ $0.652_{\pm 0.003}$
MaCoDE $0.445_{\pm 0.151}$ $0.301_{\pm 0.005}$ $0.677_{\pm 0.004}$
Dataset letter
Model $k$-anonimity(%) $\uparrow$ DCR $\uparrow$ AD $\downarrow$
CTGAN $3.520_{\pm 0.180}$ $1.583_{\pm 0.042}$ $0.089_{\pm 0.009}$
TVAE $2.657_{\pm 0.232}$ $0.779_{\pm 0.009}$ $0.274_{\pm 0.008}$
CTAB-GAN $3.343_{\pm 0.267}$ $1.438_{\pm 0.016}$ $0.100_{\pm 0.006}$
CTAB-GAN+ $3.112_{\pm 0.253}$ $1.218_{\pm 0.008}$ $0.207_{\pm 0.004}$
DistVAE $1.959_{\pm 0.265}$ $1.378_{\pm 0.006}$ $0.273_{\pm 0.005}$
TabDDPM $1.435_{\pm 0.173}$ $0.744_{\pm 0.005}$ $0.260_{\pm 0.008}$
TabMT $2.112_{\pm 0.236}$ $0.822_{\pm 0.008}$ $0.402_{\pm 0.005}$
MaCoDE $1.819_{\pm 0.089}$ $1.174_{\pm 0.011}$ $0.380_{\pm 0.007}$
Dataset loan
Model $k$-anonimity(%) $\uparrow$ DCR $\uparrow$ AD $\downarrow$
CTGAN $2.450_{\pm 0.244}$ $0.241_{\pm 0.007}$ $0.660_{\pm 0.012}$
TVAE $2.442_{\pm 0.338}$ $0.160_{\pm 0.004}$ $0.716_{\pm 0.008}$
CTAB-GAN $2.407_{\pm 0.156}$ $0.208_{\pm 0.004}$ $0.687_{\pm 0.010}$
CTAB-GAN+ $2.570_{\pm 0.159}$ $0.216_{\pm 0.015}$ $0.577_{\pm 0.005}$
DistVAE $2.348_{\pm 0.092}$ $0.238_{\pm 0.004}$ $0.683_{\pm 0.002}$
TabDDPM $0.992_{\pm 0.107}$ $0.123_{\pm 0.002}$ $0.699_{\pm 0.006}$
TabMT $1.895_{\pm 0.126}$ $0.190_{\pm 0.003}$ $0.687_{\pm 0.003}$
MaCoDE $1.942_{\pm 0.212}$ $0.127_{\pm 0.003}$ $0.697_{\pm 0.007}$
Dataset redwine
Model $k$-anonimity(%) $\uparrow$ DCR $\uparrow$ AD $\downarrow$
CTGAN $2.291_{\pm 0.411}$ $1.794_{\pm 0.032}$ $0.192_{\pm 0.034}$
TVAE $1.009_{\pm 0.325}$ $0.976_{\pm 0.016}$ $0.500_{\pm 0.041}$
CTAB-GAN $1.345_{\pm 0.140}$ $1.191_{\pm 0.027}$ $0.454_{\pm 0.027}$
CTAB-GAN+ $2.056_{\pm 0.208}$ $1.227_{\pm 0.027}$ $0.448_{\pm 0.024}$
DistVAE $3.815_{\pm 0.220}$ $1.171_{\pm 0.016}$ $0.436_{\pm 0.026}$
TabDDPM $2.697_{\pm 0.289}$ $1.330_{\pm 0.046}$ $0.480_{\pm 0.025}$
TabMT $1.071_{\pm 0.205}$ $1.236_{\pm 0.024}$ $0.420_{\pm 0.021}$
MaCoDE $1.016_{\pm 0.086}$ $0.940_{\pm 0.018}$ $0.454_{\pm 0.026}$
Dataset whitewine
Model $k$-anonimity(%) $\uparrow$ DCR $\uparrow$ AD $\downarrow$
CTGAN $2.580_{\pm 0.425}$ $1.445_{\pm 0.018}$ $0.239_{\pm 0.020}$
TVAE $1.399_{\pm 0.222}$ $1.016_{\pm 0.005}$ $0.475_{\pm 0.032}$
CTAB-GAN $1.746_{\pm 0.287}$ $1.292_{\pm 0.027}$ $0.373_{\pm 0.022}$
CTAB-GAN+ $2.116_{\pm 0.171}$ $1.145_{\pm 0.010}$ $0.385_{\pm 0.018}$
DistVAE $2.762_{\pm 0.184}$ $1.189_{\pm 0.007}$ $0.461_{\pm 0.016}$
TabDDPM $0.896_{\pm 0.064}$ $1.043_{\pm 0.009}$ $0.441_{\pm 0.009}$
TabMT $1.618_{\pm 0.303}$ $1.192_{\pm 0.013}$ $0.446_{\pm 0.014}$
MaCoDE $1.539_{\pm 0.183}$ $0.988_{\pm 0.013}$ $0.446_{\pm 0.016}$
Table 8: Privacy preservability for each dataset. The means and standard errors of the mean across 10 repeated experiments are reported. $\uparrow$ ($\downarrow$) denotes higher (lower) is better.
Sensitivity Analysis
  • Evaluation procedure: We generated missing values into the kings dataset at rates of 0.1, 0.3, 0.5, and 0.7 for each missing mechanism and proceeded to train the model. Subsequently, we evaluated the machine learning utility using metrics such as SMAPE and F1-score, along with assessing the multiple imputation performance of the model using the same methodology described earlier.
  • Result: Figure 5 and 6 show the sensitivity analysis conducted by varying the missingness rate of kings dataset across four missing data mechanisms (MCAR, MAR, MNARL, MNARQ). In Figure 5, it’s evident that MaCoDE maintains competitive performance in terms of SMAPE, even with increasing missingness rates across all missing data mechanisms. Regarding the $F_{1}$ score, MaCoDE outperforms other models at missingness rates of 0.1 and 0.3, but its performance diminishes beyond a missingness rate of 0.5 (except for the MNARQ missing data mechanism). Additionally, as shown in Figure 6, MaCoDE consistently exhibits competitive performance in the multiple imputation, regardless of the increasing missingness rate, without significant performance degradation compared to other imputation models.
    Hence, Figures 5 and 6 demonstrate that MaCoDE maintains comparable performance to other models even when trained on a dataset with missing values (i.e., incomplete dataset) without compromising the quality of the synthetic data it generates. This is a notable advantage of our proposed model over other baseline models, which struggle with training on datasets containing missing values.
Figure 5: Q2. Sensitivity analysis of machine learning utility according to missingness rate. Machine learning utility is evaluated using kings dataset under four missing mechanisms. The means and standard errors of the mean across 10 repeated experiments are reported. Error bars represent standard errors.
Figure 6: Q3. Sensitivity analysis of multiple imputations according to missingness rate. Multiple imputation is evaluated using kings dataset under four missing mechanisms. The means and standard errors of the mean across 10 repeated experiments are reported. Error bars represent standard errors.
Detailed Experimental Results
Q1. Synthetic Data Quality
abalone
Statistical fidelity Machine learning utility
Model KL $\downarrow$ GoF $\downarrow$ MMD $\downarrow$ WD $\downarrow$ SMAPE $\downarrow$ $F_{1}$uparrow Model $\uparrow$ Feature $\uparrow$
Baseline .$006_{\pm.000}.$$023_{\pm.001}.$$001_{\pm.000}.$$152_{\pm.010}.$$031_{\pm.000}.$$229_{\pm.003}.$$790_{\pm.057}.$$981_{\pm.006}.$
CTGAN .$114_{\pm.011}.$$289_{\pm.058}.$$109_{\pm.010}.$ $1.174_{\pm.069}$$101_{\pm.003}.$$166_{\pm.004}.$$625_{\pm.087}.$$283_{\pm.095}.$
TVAE .$064_{\pm.003}.$$121_{\pm.004}.$$016_{\pm.001}.$$387_{\pm.007}.$$120_{\pm.005}.$$247_{\pm.003}.$$152_{\pm.146}.$$417_{\pm.105}.$
CTAB-GAN .$208_{\pm.012}.$$403_{\pm.054}.$$096_{\pm.013}.$ $1.161_{\pm.120}$$109_{\pm.005}.$$162_{\pm.005}.$$274_{\pm.161}.$$395_{\pm.158}.$
CTAB-GAN+ .$035_{\pm.008}.$$091_{\pm.020}.$$022_{\pm.006}.$$485_{\pm.094}.$$068_{\pm.006}.$$189_{\pm.013}.$$805_{\pm.147}.$$771_{\pm.100}.$
DistVAE .$010_{\pm.001}.$$043_{\pm.003}.$$005_{\pm.000}.$$263_{\pm.007}.$$061_{\pm.001}.$$216_{\pm.002}.$$825_{\pm.050}.$$288_{\pm.034}.$
TabDDPM .$119_{\pm.030}.$$905_{\pm.577}.$$005_{\pm.002}.$ $4.592_{\pm 1.055}$$033_{\pm.000}.$$222_{\pm.002}.$$747_{\pm.059}.$$957_{\pm.008}.$
TabMT .$003_{\pm.000}.$$022_{\pm.001}.$$033_{\pm.001}.$$949_{\pm.033}.$$081_{\pm.001}.$$198_{\pm.003}.$$744_{\pm.050}.$$852_{\pm.022}.$
MaCoDE .$017_{\pm.002}.$$064_{\pm.006}.$$005_{\pm.001}.$$150_{\pm.009}.$$039_{\pm.000}.$$217_{\pm.003}.$$744_{\pm.049}.$$952_{\pm.010}.$
banknote
Statistical fidelity Machine learning utility
Model KL $\downarrow$ GoF $\downarrow$ MMD $\downarrow$ WD $\downarrow$ SMAPE $\downarrow$ $F_{1}$uparrow Model $\uparrow$ Feature $\uparrow$
Baseline .$016_{\pm.001}.$$040_{\pm.003}.$$002_{\pm.000}.$$059_{\pm.004}.$$232_{\pm.002}.$$957_{\pm.002}.$$903_{\pm.027}.$ $1.000_{\pm.000}$
CTGAN .$146_{\pm.013}.$$187_{\pm.013}.$$095_{\pm.009}.$ $1.260_{\pm.119}$$646_{\pm.009}.$$745_{\pm.020}.$$-.321_{\pm.143}- .$$720_{\pm.085}.$
TVAE .$047_{\pm.009}.$$084_{\pm.008}.$$019_{\pm.004}.$$260_{\pm.023}.$$436_{\pm.006}.$$879_{\pm.006}.$$616_{\pm.079}.$$940_{\pm.031}.$
CTAB-GAN .$030_{\pm.003}.$$066_{\pm.003}.$$021_{\pm.001}.$$415_{\pm.018}.$$519_{\pm.004}.$$826_{\pm.011}.$$207_{\pm.166}.$$880_{\pm.033}.$
CTAB-GAN+ .$176_{\pm.076}.$$137_{\pm.042}.$$072_{\pm.038}.$$847_{\pm.227}.$$670_{\pm.031}.$$634_{\pm.059}.$$-.388_{\pm.274}- .$$280_{\pm.605}.$
DistVAE .$075_{\pm.002}.$$068_{\pm.001}.$$031_{\pm.001}.$$563_{\pm.011}.$$651_{\pm.005}.$$763_{\pm.009}.$$-.513_{\pm.081}- .$ $1.000_{\pm.000}$
TabDDPM .$052_{\pm.009}.$$054_{\pm.004}.$$006_{\pm.001}.$$364_{\pm.059}.$$315_{\pm.004}.$$942_{\pm.003}.$$839_{\pm.033}.$ $1.000_{\pm.000}$
TabMT .$012_{\pm.001}.$$033_{\pm.001}.$$012_{\pm.001}.$$301_{\pm.008}.$$495_{\pm.005}.$$883_{\pm.007}.$$468_{\pm.126}.$$960_{\pm.027}.$
MaCoDE .$031_{\pm.003}.$$074_{\pm.005}.$$009_{\pm.001}.$$148_{\pm.011}.$$374_{\pm.005}.$$918_{\pm.005}.$$652_{\pm.071}.$ $1.000_{\pm.000}$
breast
Statistical fidelity Machine learning utility
Model KL $\downarrow$ GoF $\downarrow$ MMD $\downarrow$ WD $\downarrow$ SMAPE $\downarrow$ $F_{1}$uparrow Model $\uparrow$ Feature $\uparrow$
Baseline .$081_{\pm.005}.$$080_{\pm.003}.$$007_{\pm.001}.$ $5.511_{\pm.116}$$056_{\pm.001}.$$946_{\pm.002}.$$662_{\pm.106}.$$865_{\pm.015}.$
CTGAN .$419_{\pm.029}.$$345_{\pm.013}.$$242_{\pm.008}.$ $35.684_{\pm 1.061}$$237_{\pm.006}.$$645_{\pm.046}.$$057_{\pm.167}.$$-.163_{\pm.045}- .$
TVAE .$049_{\pm.002}.$$087_{\pm.002}.$$010_{\pm.001}.$ $6.519_{\pm.079}$$090_{\pm.001}.$$938_{\pm.004}.$$687_{\pm.071}.$$895_{\pm.012}.$
CTAB-GAN .$218_{\pm.010}.$$186_{\pm.005}.$$092_{\pm.005}.$ $16.515_{\pm.534}$$151_{\pm.005}.$$877_{\pm.010}.$$350_{\pm.129}.$$558_{\pm.034}.$
CTAB-GAN+ .$564_{\pm.255}.$$356_{\pm.092}.$$229_{\pm.090}.$ $25.009_{\pm 9.867}$$242_{\pm.046}.$$696_{\pm.138}.$$265_{\pm.525}.$$187_{\pm.275}.$
DistVAE .$071_{\pm.002}.$$090_{\pm.001}.$$033_{\pm.001}.$ $1.407_{\pm.107}$$123_{\pm.002}.$$889_{\pm.006}.$$274_{\pm.071}.$$750_{\pm.022}.$
TabDDPM $3.110_{\pm.025}$$587_{\pm.006}.$$306_{\pm.001}.$ $262.226_{\pm 6.849}$$345_{\pm.008}.$$638_{\pm.016}.$$-.260_{\pm.151}- .$$364_{\pm.041}.$
TabMT .$032_{\pm.002}.$$060_{\pm.002}.$$028_{\pm.001}.$ $12.405_{\pm.123}$$128_{\pm.002}.$$911_{\pm.003}.$$503_{\pm.131}.$$756_{\pm.017}.$
MaCoDE .$036_{\pm.003}.$$079_{\pm.006}.$$012_{\pm.001}.$ $8.634_{\pm.122}$$098_{\pm.002}.$$921_{\pm.004}.$$332_{\pm.167}.$$807_{\pm.010}.$
concrete
Statistical fidelity Machine learning utility
Model KL $\downarrow$ GoF $\downarrow$ MMD $\downarrow$ WD $\downarrow$ SMAPE $\downarrow$ $F_{1}$uparrow Model $\uparrow$ Feature $\uparrow$
Baseline .$026_{\pm.002}.$$055_{\pm.002}.$$003_{\pm.000}.$$433_{\pm.006}.$$121_{\pm.002}.$$466_{\pm.009}.$$912_{\pm.017}.$$950_{\pm.009}.$
CTGAN .$384_{\pm.027}.$$830_{\pm.129}.$$134_{\pm.012}.$ $4.443_{\pm.289}$$334_{\pm.004}.$$091_{\pm.020}.$$070_{\pm.151}.$$114_{\pm.138}.$
TVAE .$125_{\pm.007}.$$219_{\pm.007}.$$017_{\pm.002}.$ $1.057_{\pm.031}$$276_{\pm.003}.$$398_{\pm.010}.$$030_{\pm.161}.$$290_{\pm.045}.$
CTAB-GAN .$204_{\pm.015}.$$226_{\pm.010}.$$051_{\pm.006}.$ $2.383_{\pm.098}$$290_{\pm.004}.$$325_{\pm.011}.$$035_{\pm.061}.$$-.260_{\pm.097}- .$
CTAB-GAN+ .$263_{\pm.054}.$$278_{\pm.027}.$$066_{\pm.023}.$ $2.472_{\pm.380}$$285_{\pm.012}.$$331_{\pm.045}.$$-.009_{\pm.175}- .$$017_{\pm.344}.$
DistVAE .$186_{\pm.026}.$$159_{\pm.002}.$$024_{\pm.001}.$ $1.808_{\pm.011}$$289_{\pm.003}.$$374_{\pm.007}.$$-.160_{\pm.092}- .$$107_{\pm.104}.$
TabDDPM .$036_{\pm.001}.$$087_{\pm.007}.$$001_{\pm.000}.$$532_{\pm.031}.$$156_{\pm.004}.$$446_{\pm.007}.$$537_{\pm.049}.$$938_{\pm.019}.$
TabMT .$031_{\pm.001}.$$115_{\pm.002}.$$019_{\pm.001}.$ $1.810_{\pm.034}$$284_{\pm.003}.$$355_{\pm.010}.$$392_{\pm.060}.$$121_{\pm.086}.$
MaCoDE .$089_{\pm.019}.$$118_{\pm.008}.$$006_{\pm.001}.$$684_{\pm.025}.$$229_{\pm.002}.$$405_{\pm.009}.$$439_{\pm.081}.$$919_{\pm.021}.$
covtype
Statistical fidelity Machine learning utility
Model KL $\downarrow$ GoF $\downarrow$ MMD $\downarrow$ WD $\downarrow$ SMAPE $\downarrow$ $F_{1}$uparrow Model $\uparrow$ Feature $\uparrow$
Baseline .$000_{\pm.000}.$$002_{\pm.000}.$$000_{\pm.000}.$$454_{\pm.002}.$$067_{\pm.000}.$$804_{\pm.000}.$ $1.000_{\pm.000}$ $1.000_{\pm.000}$
CTGAN .$105_{\pm.007}.$$493_{\pm.063}.$$017_{\pm.002}.$ $1.338_{\pm.053}$$154_{\pm.001}.$$495_{\pm.023}.$$320_{\pm.098}.$$878_{\pm.011}.$
TVAE .$048_{\pm.004}.$$078_{\pm.004}.$$013_{\pm.002}.$$950_{\pm.021}.$$151_{\pm.001}.$$634_{\pm.008}.$$420_{\pm.055}.$$930_{\pm.019}.$
CTAB-GAN .$020_{\pm.002}.$$064_{\pm.005}.$$005_{\pm.000}.$$892_{\pm.019}.$$147_{\pm.001}.$$574_{\pm.005}.$$250_{\pm.067}.$$922_{\pm.010}.$
CTAB-GAN+ .$002_{\pm.000}.$$020_{\pm.003}.$$001_{\pm.000}.$$601_{\pm.007}.$$122_{\pm.001}.$$652_{\pm.003}.$$300_{\pm.000}.$$958_{\pm.006}.$
DistVAE .$007_{\pm.000}.$$028_{\pm.001}.$$005_{\pm.000}.$ $1.228_{\pm.010}$$182_{\pm.001}.$$626_{\pm.002}.$$-.080_{\pm.033}- .$$895_{\pm.016}.$
TabDDPM .$003_{\pm.000}.$$011_{\pm.000}.$$001_{\pm.000}.$$556_{\pm.006}.$$115_{\pm.000}.$$665_{\pm.001}.$$300_{\pm.000}.$$964_{\pm.000}.$
TabMT .$001_{\pm.000}.$$009_{\pm.000}.$$001_{\pm.000}.$$621_{\pm.005}.$$119_{\pm.000}.$$701_{\pm.002}.$$700_{\pm.000}.$$960_{\pm.002}.$
MaCoDE .$001_{\pm.000}.$$026_{\pm.001}.$$001_{\pm.000}.$$509_{\pm.002}.$$098_{\pm.000}.$$742_{\pm.001}.$$900_{\pm.000}.$$956_{\pm.002}.$
Table 9: Q1: Statistical fidelity and machine learning utility for each dataset. The means and the standard errors of the mean across 10 repeated experiments are reported. ‘Baseline’ refers to the result obtained using half of the real training dataset. $\uparrow$ ($\downarrow$) denotes higher (lower) is better.
kings
Statistical fidelity Machine learning utility
Model KL $\downarrow$ GoF $\downarrow$ MMD $\downarrow$ WD $\downarrow$ SMAPE $\downarrow$ $F_{1}$uparrow Model $\uparrow$ Feature $\uparrow$
Baseline .$002_{\pm.000}.$$008_{\pm.000}.$$000_{\pm.000}.$$414_{\pm.014}.$$088_{\pm.001}.$$605_{\pm.003}.$ $1.000_{\pm.000}$$994_{\pm.001}.$
CTGAN .$141_{\pm.009}.$$753_{\pm.083}.$$036_{\pm.004}.$ $2.702_{\pm.347}$$245_{\pm.003}.$$437_{\pm.010}.$$880_{\pm.025}.$$881_{\pm.008}.$
TVAE .$032_{\pm.001}.$$097_{\pm.002}.$$004_{\pm.001}.$$942_{\pm.038}.$$204_{\pm.000}.$$573_{\pm.003}.$$940_{\pm.016}.$$894_{\pm.008}.$
CTAB-GAN .$056_{\pm.004}.$$421_{\pm.223}.$$017_{\pm.001}.$ $1.388_{\pm.114}$$222_{\pm.002}.$$466_{\pm.010}.$$947_{\pm.016}.$$898_{\pm.008}.$
CTAB-GAN+ .$037_{\pm.030}.$$100_{\pm.048}.$$039_{\pm.037}.$ $2.107_{\pm 1.594}$$181_{\pm.007}.$$526_{\pm.022}.$$937_{\pm.055}.$$956_{\pm.022}.$
DistVAE .$159_{\pm.017}.$$085_{\pm.002}.$$005_{\pm.000}.$ $1.039_{\pm.035}$$236_{\pm.001}.$$541_{\pm.004}.$$930_{\pm.030}.$$907_{\pm.003}.$
TabDDPM .$096_{\pm.017}.$ $1.621_{\pm.684}$$001_{\pm.000}.$ $27.567_{\pm 6.882}$$083_{\pm.001}.$$549_{\pm.007}.$$900_{\pm.038}.$$978_{\pm.005}.$
TabMT .$001_{\pm.000}.$$010_{\pm.000}.$$001_{\pm.000}.$$741_{\pm.011}.$$154_{\pm.001}.$$545_{\pm.005}.$$940_{\pm.016}.$$964_{\pm.002}.$
MaCoDE .$011_{\pm.001}.$$046_{\pm.002}.$$009_{\pm.002}.$$998_{\pm.089}.$$179_{\pm.001}.$$580_{\pm.004}.$$920_{\pm.025}.$$975_{\pm.002}.$
letter
Statistical fidelity Machine learning utility
Model KL $\downarrow$ GoF $\downarrow$ MMD $\downarrow$ WD $\downarrow$ SMAPE $\downarrow$ $F_{1}$uparrow Model $\uparrow$ Feature $\uparrow$
Baseline .$001_{\pm.000}.$$010_{\pm.000}.$$000_{\pm.000}.$$921_{\pm.002}.$$059_{\pm.000}.$$821_{\pm.002}.$ $1.000_{\pm.000}$$990_{\pm.001}.$
CTGAN .$171_{\pm.004}.$$294_{\pm.004}.$$022_{\pm.002}.$ $4.743_{\pm.182}$$169_{\pm.003}.$$169_{\pm.010}.$$120_{\pm.135}.$$555_{\pm.055}.$
TVAE .$040_{\pm.002}.$$064_{\pm.002}.$$005_{\pm.000}.$ $1.604_{\pm.013}$$106_{\pm.001}.$$496_{\pm.011}.$$550_{\pm.056}.$$944_{\pm.005}.$
CTAB-GAN .$113_{\pm.005}.$$121_{\pm.008}.$$013_{\pm.001}.$ $3.921_{\pm.059}$$153_{\pm.002}.$$197_{\pm.010}.$$130_{\pm.087}.$$729_{\pm.030}.$
CTAB-GAN+ .$051_{\pm.012}.$$073_{\pm.015}.$$005_{\pm.001}.$ $2.782_{\pm.112}$$123_{\pm.001}.$$471_{\pm.021}.$$440_{\pm.126}.$$899_{\pm.023}.$
DistVAE .$021_{\pm.001}.$$044_{\pm.001}.$$004_{\pm.000}.$ $2.526_{\pm.015}$$120_{\pm.000}.$$577_{\pm.003}.$$110_{\pm.046}.$$854_{\pm.009}.$
TabDDPM .$009_{\pm.000}.$$031_{\pm.001}.$$002_{\pm.000}.$ $1.371_{\pm.023}$$075_{\pm.000}.$$448_{\pm.005}.$$500_{\pm.001}.$$531_{\pm.008}.$
TabMT .$002_{\pm.000}.$$010_{\pm.000}.$$001_{\pm.000}.$ $1.579_{\pm.009}$$083_{\pm.000}.$$745_{\pm.003}.$$900_{\pm.001}.$$976_{\pm.003}.$
MaCoDE .$095_{\pm.002}.$$144_{\pm.003}.$$017_{\pm.001}.$ $2.135_{\pm.035}$$106_{\pm.001}.$$689_{\pm.003}.$$850_{\pm.017}.$$965_{\pm.003}.$
loan
Statistical fidelity Machine learning utility
Model KL $\downarrow$ GoF $\downarrow$ MMD $\downarrow$ WD $\downarrow$ SMAPE $\downarrow$ $F_{1}$uparrow Model $\uparrow$ Feature $\uparrow$
Baseline .$002_{\pm.000}.$$011_{\pm.000}.$$001_{\pm.000}.$$071_{\pm.001}.$$267_{\pm.001}.$$930_{\pm.001}.$$907_{\pm.037}.$$965_{\pm.008}.$
CTGAN .$079_{\pm.009}.$$177_{\pm.014}.$$056_{\pm.006}.$$963_{\pm.070}.$$368_{\pm.003}.$$886_{\pm.004}.$$272_{\pm.079}.$$682_{\pm.045}.$
TVAE .$121_{\pm.004}.$$170_{\pm.003}.$$023_{\pm.000}.$$751_{\pm.028}.$$339_{\pm.002}.$$850_{\pm.007}.$$685_{\pm.102}.$$916_{\pm.017}.$
CTAB-GAN .$037_{\pm.004}.$$087_{\pm.004}.$$023_{\pm.003}.$$502_{\pm.032}.$$333_{\pm.004}.$$904_{\pm.003}.$$440_{\pm.097}.$$802_{\pm.014}.$
CTAB-GAN+ .$050_{\pm.028}.$$084_{\pm.009}.$$039_{\pm.024}.$$505_{\pm.251}.$$304_{\pm.020}.$$892_{\pm.011}.$$144_{\pm.313}.$$810_{\pm.043}.$
DistVAE .$011_{\pm.001}.$$058_{\pm.002}.$$030_{\pm.000}.$$673_{\pm.007}.$$361_{\pm.001}.$$903_{\pm.003}.$$517_{\pm.069}.$$781_{\pm.016}.$
TabDDPM .$004_{\pm.000}.$$010_{\pm.000}.$$001_{\pm.000}.$$075_{\pm.007}.$$281_{\pm.001}.$$925_{\pm.001}.$$677_{\pm.039}.$$978_{\pm.006}.$
TabMT .$002_{\pm.000}.$$013_{\pm.001}.$$003_{\pm.000}.$$192_{\pm.008}.$$307_{\pm.001}.$$913_{\pm.002}.$$623_{\pm.043}.$$891_{\pm.012}.$
MaCoDE .$031_{\pm.004}.$$047_{\pm.003}.$$007_{\pm.001}.$$126_{\pm.008}.$$273_{\pm.001}.$$919_{\pm.002}.$$726_{\pm.033}.$$896_{\pm.021}.$
redwine
Statistical fidelity Machine learning utility
Model KL $\downarrow$ GoF $\downarrow$ MMD $\downarrow$ WD $\downarrow$ SMAPE $\downarrow$ $F_{1}$uparrow Model $\uparrow$ Feature $\uparrow$
Baseline .$025_{\pm.004}.$$042_{\pm.002}.$$003_{\pm.000}.$ $1.180_{\pm.027}$$080_{\pm.001}.$$576_{\pm.005}.$$802_{\pm.072}.$$905_{\pm.014}.$
CTGAN .$463_{\pm.017}.$$914_{\pm.178}.$$164_{\pm.007}.$ $8.166_{\pm.528}$$177_{\pm.005}.$$228_{\pm.031}.$$116_{\pm.144}.$$071_{\pm.119}.$
TVAE .$057_{\pm.002}.$$115_{\pm.002}.$$012_{\pm.001}.$ $1.842_{\pm.042}$$103_{\pm.001}.$$569_{\pm.005}.$$238_{\pm.117}.$$684_{\pm.043}.$
CAB-GAN .$111_{\pm.012}.$$143_{\pm.008}.$$042_{\pm.006}.$ $3.075_{\pm.150}$$139_{\pm.003}.$$506_{\pm.007}.$$010_{\pm.109}.$$445_{\pm.056}.$
CTAB-GAN+ .$118_{\pm.025}.$$157_{\pm.031}.$$044_{\pm.014}.$ $2.969_{\pm.312}$$152_{\pm.012}.$$488_{\pm.049}.$$-.145_{\pm.413}- .$$510_{\pm.343}.$
DistVAE .$035_{\pm.001}.$$064_{\pm.001}.$$015_{\pm.000}.$ $2.423_{\pm.018}$$134_{\pm.001}.$$533_{\pm.005}.$$-.108_{\pm.091}- .$$776_{\pm.030}.$
TabDDPM $2.678_{\pm.108}$$652_{\pm.050}.$$214_{\pm.009}.$ $86.984_{\pm 4.091}$$108_{\pm.002}.$$481_{\pm.010}.$$412_{\pm.122}.$$372_{\pm.066}.$
TabMT .$023_{\pm.002}.$$052_{\pm.001}.$$015_{\pm.000}.$ $2.534_{\pm.024}$$127_{\pm.002}.$$507_{\pm.009}.$$010_{\pm.112}.$$567_{\pm.057}.$
MaCoDE .$022_{\pm.001}.$$071_{\pm.002}.$$005_{\pm.000}.$ $1.526_{\pm.016}$$105_{\pm.001}.$$524_{\pm.007}.$$197_{\pm.092}.$$870_{\pm.032}.$
whitewine
Statistical fidelity Machine learning utility
Model KL $\downarrow$ GoF $\downarrow$ MMD $\downarrow$ WD $\downarrow$ SMAPE $\downarrow$ $F_{1}$uparrow Model $\uparrow$ Feature $\uparrow$
Baseline .$006_{\pm.001}.$$022_{\pm.001}.$$001_{\pm.000}.$$993_{\pm.008}.$$065_{\pm.000}.$$523_{\pm.004}.$$897_{\pm.027}.$$913_{\pm.012}.$
CTGAN .$186_{\pm.008}.$ $1.330_{\pm.159}$$068_{\pm.005}.$ $3.881_{\pm.153}$$123_{\pm.002}.$$248_{\pm.019}.$$-.060_{\pm.064}- .$$148_{\pm.083}.$
TVAE .$078_{\pm.005}.$$152_{\pm.005}.$$040_{\pm.004}.$ $1.996_{\pm.064}$$100_{\pm.001}.$$492_{\pm.003}.$$545_{\pm.127}.$$561_{\pm.066}.$
CTAB-GAN .$166_{\pm.016}.$$241_{\pm.019}.$$080_{\pm.011}.$ $3.016_{\pm.147}$$117_{\pm.002}.$$407_{\pm.010}.$$-.010_{\pm.089}- .$$311_{\pm.090}.$
CTAB-GAN+ .$064_{\pm.022}.$$139_{\pm.022}.$$029_{\pm.012}.$ $1.929_{\pm.265}$$109_{\pm.004}.$$424_{\pm.021}.$$-.080_{\pm.326}- .$$620_{\pm.189}.$
DistVAE .$018_{\pm.001}.$$059_{\pm.001}.$$009_{\pm.001}.$ $1.787_{\pm.017}$$106_{\pm.001}.$$459_{\pm.003}.$$150_{\pm.095}.$$595_{\pm.026}.$
TabDDPM .$675_{\pm.041}.$$161_{\pm.011}.$$022_{\pm.002}.$ $4.292_{\pm 2.951}$$079_{\pm.001}.$$439_{\pm.008}.$$535_{\pm.070}.$$685_{\pm.056}.$
TabMT .$005_{\pm.001}.$$031_{\pm.000}.$$007_{\pm.000}.$ $1.861_{\pm.010}$$102_{\pm.001}.$$462_{\pm.004}.$$-.001_{\pm.068}- .$$562_{\pm.035}.$
MaCoDE .$010_{\pm.001}.$$049_{\pm.002}.$$002_{\pm.000}.$ $1.238_{\pm.008}$$087_{\pm.000}.$$465_{\pm.003}.$$240_{\pm.070}.$$903_{\pm.013}.$
Table 10: Q1: Statistical fidelity and machine learning utility for each dataset. The means and the standard errors of the mean across 10 repeated experiments are reported. ‘Baseline’ refers to the result obtained using half of the real training dataset. $\uparrow$ ($\downarrow$) denotes higher (lower) is better.
Q1. Visualization of Marginal Histogram
(a) abalone
(b) banknote
(c) breast
(d) concreate
(e) covtype
Figure 7: Histograms of observed dataset and synthetic dataset, generated by MaCoDE.
(a) kings
(b) letter
(c) loan
(d) redwine
(e) whitewine
Figure 8: Histograms of observed dataset and synthetic dataset, generated by MaCoDE.
Q2: Synthetic Data Quality in Scenarios with Incomplete Training Dataset
MCAR
Dataset SMAPE $\downarrow$ $F_{1}$uparrow
abalone $0.042_{\pm 0.000}$ $0.216_{\pm 0.003}$
banknote $0.417_{\pm 0.004}$ $0.879_{\pm 0.006}$
breast $0.111_{\pm 0.001}$ $0.909_{\pm 0.006}$
concrete $0.240_{\pm 0.002}$ $0.373_{\pm 0.010}$
covtype $0.103_{\pm 0.000}$ $0.721_{\pm 0.001}$
kings $0.180_{\pm 0.003}$ $0.565_{\pm 0.004}$
letter $0.107_{\pm 0.000}$ $0.674_{\pm 0.004}$
loan $0.278_{\pm 0.003}$ $0.915_{\pm 0.002}$
redwine $0.112_{\pm 0.001}$ $0.515_{\pm 0.009}$
whitewine $0.091_{\pm 0.000}$ $0.461_{\pm 0.005}$
MAR
Dataset SMAPE $\downarrow$ $F_{1}$uparrow
abalone $0.041_{\pm 0.000}$ $0.211_{\pm 0.003}$
banknote $0.410_{\pm 0.005}$ $0.891_{\pm 0.005}$
breast $0.107_{\pm 0.002}$ $0.916_{\pm 0.005}$
concrete $0.244_{\pm 0.003}$ $0.383_{\pm 0.007}$
covtype $0.102_{\pm 0.000}$ $0.723_{\pm 0.001}$
kings $0.183_{\pm 0.001}$ $0.561_{\pm 0.005}$
letter $0.107_{\pm 0.000}$ $0.682_{\pm 0.004}$
loan $0.278_{\pm 0.001}$ $0.914_{\pm 0.002}$
redwine $0.110_{\pm 0.001}$ $0.518_{\pm 0.005}$
whitewine $0.090_{\pm 0.001}$ $0.465_{\pm 0.006}$
MNARL
Dataset SMAPE $\downarrow$ $F_{1}$uparrow
abalone $0.042_{\pm 0.000}$ $0.212_{\pm 0.004}$
banknote $0.418_{\pm 0.007}$ $0.883_{\pm 0.006}$
breast $0.112_{\pm 0.002}$ $0.908_{\pm 0.006}$
concrete $0.243_{\pm 0.003}$ $0.374_{\pm 0.011}$
covtype $0.104_{\pm 0.000}$ $0.718_{\pm 0.001}$
kings $0.182_{\pm 0.004}$ $0.560_{\pm 0.009}$
letter $0.108_{\pm 0.001}$ $0.676_{\pm 0.003}$
loan $0.277_{\pm 0.001}$ $0.915_{\pm 0.002}$
redwine $0.112_{\pm 0.001}$ $0.511_{\pm 0.009}$
whitewine $0.091_{\pm 0.001}$ $0.464_{\pm 0.006}$
MNARQ
Dataset SMAPE $\downarrow$ $F_{1}$uparrow
abalone $0.041_{\pm 0.000}$ $0.211_{\pm 0.003}$
banknote $0.399_{\pm 0.003}$ $0.899_{\pm 0.009}$
breast $0.105_{\pm 0.001}$ $0.918_{\pm 0.003}$
concrete $0.239_{\pm 0.003}$ $0.384_{\pm 0.009}$
covtype $0.101_{\pm 0.000}$ $0.730_{\pm 0.001}$
kings $0.179_{\pm 0.004}$ $0.576_{\pm 0.003}$
letter $0.107_{\pm 0.001}$ $0.676_{\pm 0.004}$
loan $0.275_{\pm 0.001}$ $0.916_{\pm 0.001}$
redwine $0.108_{\pm 0.001}$ $0.527_{\pm 0.006}$
whitewine $0.089_{\pm 0.001}$ $0.460_{\pm 0.004}$
Table 11: Q2: Machine learning utility for each dataset under MCAR, MAR, MNARL, and MNARQ at 0.3 missingness. The means and standard errors of the mean 10 repeated experiments are reported. $\uparrow$ ($\downarrow$) denotes higher (lower) is better.
Q3: Multiple Imputation Performance
MCAR
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.014_{\pm 0.001}$ $0.795_{\pm 0.022}$ $0.040_{\pm 0.002}$
GAIN $0.024_{\pm 0.002}$ $0.591_{\pm 0.037}$ $0.040_{\pm 0.002}$
missMDA $0.018_{\pm 0.001}$ $0.644_{\pm 0.015}$ $0.045_{\pm 0.002}$
VAEAC $0.010_{\pm 0.001}$ $0.874_{\pm 0.022}$ $0.041_{\pm 0.002}$
MIWAE $0.008_{\pm 0.000}$ $0.936_{\pm 0.014}$ $0.045_{\pm 0.002}$
notMIWAE $0.008_{\pm 0.000}$ $0.910_{\pm 0.013}$ $0.044_{\pm 0.002}$
EGC $0.005_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.060_{\pm 0.002}$
MaCoDE(MCAR) $0.008_{\pm 0.000}$ $0.950_{\pm 0.011}$ $0.053_{\pm 0.002}$
MAR
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.010_{\pm 0.001}$ $0.845_{\pm 0.019}$ $0.040_{\pm 0.002}$
GAIN $0.019_{\pm 0.002}$ $0.633_{\pm 0.033}$ $0.040_{\pm 0.002}$
missMDA $0.015_{\pm 0.001}$ $0.700_{\pm 0.022}$ $0.043_{\pm 0.002}$
VAEAC $0.008_{\pm 0.001}$ $0.905_{\pm 0.016}$ $0.040_{\pm 0.002}$
MIWAE $0.006_{\pm 0.000}$ $0.952_{\pm 0.012}$ $0.043_{\pm 0.002}$
notMIWAE $0.006_{\pm 0.000}$ $0.949_{\pm 0.012}$ $0.042_{\pm 0.002}$
EGC $0.006_{\pm 0.000}$ $0.996_{\pm 0.004}$ $0.058_{\pm 0.002}$
MaCoDE(MAR) $0.006_{\pm 0.000}$ $0.963_{\pm 0.009}$ $0.051_{\pm 0.003}$
MNARL
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.012_{\pm 0.001}$ $0.790_{\pm 0.026}$ $0.037_{\pm 0.002}$
GAIN $0.026_{\pm 0.002}$ $0.538_{\pm 0.033}$ $0.040_{\pm 0.002}$
missMDA $0.020_{\pm 0.001}$ $0.592_{\pm 0.025}$ $0.045_{\pm 0.002}$
VAEAC $0.011_{\pm 0.001}$ $0.851_{\pm 0.023}$ $0.041_{\pm 0.002}$
MIWAE $0.009_{\pm 0.001}$ $0.920_{\pm 0.015}$ $0.045_{\pm 0.002}$
notMIWAE $0.009_{\pm 0.001}$ $0.908_{\pm 0.012}$ $0.044_{\pm 0.002}$
EGC $0.007_{\pm 0.000}$ $0.994_{\pm 0.004}$ $0.061_{\pm 0.002}$
MaCoDE(MNARL) $0.008_{\pm 0.000}$ $0.948_{\pm 0.013}$ $0.052_{\pm 0.003}$
MNARQ
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.007_{\pm 0.001}$ $0.880_{\pm 0.018}$ $0.040_{\pm 0.002}$
GAIN $0.011_{\pm 0.001}$ $0.819_{\pm 0.020}$ $0.040_{\pm 0.002}$
missMDA $0.011_{\pm 0.001}$ $0.785_{\pm 0.016}$ $0.043_{\pm 0.002}$
VAEAC $0.006_{\pm 0.001}$ $0.909_{\pm 0.020}$ $0.040_{\pm 0.002}$
MIWAE $0.005_{\pm 0.000}$ $0.969_{\pm 0.009}$ $0.043_{\pm 0.002}$
notMIWAE $0.005_{\pm 0.000}$ $0.958_{\pm 0.010}$ $0.043_{\pm 0.002}$
EGC $0.004_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.054_{\pm 0.002}$
MaCoDE(MNARQ) $0.005_{\pm 0.000}$ $0.944_{\pm 0.012}$ $0.051_{\pm 0.003}$
Table 12: Q3: Multiple imputation under MCAR, MAR, MNARL, and MNARQ at 0.3 missingness. The means and standard errors of the mean across 5 datasets and 10 repeated experiments are reported. $\downarrow$ denotes the lower is better.
Dataset abalone
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.005_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.030_{\pm 0.000}$
GAIN $0.009_{\pm 0.001}$ $0.829_{\pm 0.042}$ $0.030_{\pm 0.000}$
missMDA $0.014_{\pm 0.000}$ $0.686_{\pm 0.036}$ $0.034_{\pm 0.001}$
VAEAC $0.005_{\pm 0.002}$ $0.943_{\pm 0.074}$ $0.030_{\pm 0.000}$
MIWAE $0.006_{\pm 0.001}$ $0.971_{\pm 0.019}$ $0.034_{\pm 0.001}$
not-MIWAE $0.009_{\pm 0.001}$ $0.886_{\pm 0.019}$ $0.037_{\pm 0.001}$
EGC $0.004_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.040_{\pm 0.000}$
MaCoDE $0.007_{\pm 0.001}$ $0.914_{\pm 0.023}$ $0.036_{\pm 0.002}$
Dataset banknote
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.018_{\pm 0.001}$ $0.725_{\pm 0.025}$ $0.053_{\pm 0.000}$
GAIN $0.009_{\pm 0.001}$ $0.829_{\pm 0.042}$ $0.030_{\pm 0.000}$
missMDA $0.018_{\pm 0.000}$ $0.625_{\pm 0.042}$ $0.056_{\pm 0.000}$
VAEAC $0.009_{\pm 0.002}$ $1.000_{\pm 0.000}$ $0.054_{\pm 0.000}$
MIWAE $0.006_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.055_{\pm 0.000}$
not-MIWAE $0.006_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.055_{\pm 0.000}$
EGC $0.004_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.066_{\pm 0.001}$
MaCoDE $0.006_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.054_{\pm 0.000}$
Dataset breast
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.007_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.080_{\pm 0.000}$
GAIN $0.012_{\pm 0.000}$ $0.990_{\pm 0.005}$ $0.080_{\pm 0.000}$
missMDA $0.025_{\pm 0.000}$ $0.883_{\pm 0.013}$ $0.086_{\pm 0.000}$
VAEAC $0.008_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.082_{\pm 0.000}$
MIWAE $0.005_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.084_{\pm 0.000}$
not-MIWAE $0.006_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.082_{\pm 0.000}$
EGC $0.005_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.084_{\pm 0.000}$
MaCoDE $0.009_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.083_{\pm 0.000}$
Dataset redwine
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.020_{\pm 0.000}$ $0.764_{\pm 0.020}$ $0.048_{\pm 0.000}$
GAIN $0.024_{\pm 0.001}$ $0.627_{\pm 0.050}$ $0.048_{\pm 0.000}$
missMDA $0.026_{\pm 0.000}$ $0.618_{\pm 0.012}$ $0.055_{\pm 0.000}$
VAEAC $0.014_{\pm 0.001}$ $0.855_{\pm 0.028}$ $0.050_{\pm 0.001}$
MIWAE $0.009_{\pm 0.001}$ $0.945_{\pm 0.020}$ $0.054_{\pm 0.000}$
not-MIWAE $0.010_{\pm 0.000}$ $0.909_{\pm 0.000}$ $0.053_{\pm 0.000}$
EGC $0.007_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.070_{\pm 0.000}$
MaCoDE $0.008_{\pm 0.001}$ $0.973_{\pm 0.014}$ $0.056_{\pm 0.000}$
Dataset whitewine
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.011_{\pm 0.000}$ $0.691_{\pm 0.031}$ $0.028_{\pm 0.000}$
GAIN $0.026_{\pm 0.003}$ $0.482_{\pm 0.036}$ $0.028_{\pm 0.000}$
missMDA $0.015_{\pm 0.000}$ $0.645_{\pm 0.016}$ $0.035_{\pm 0.000}$
VAEAC $0.010_{\pm 0.001}$ $0.700_{\pm 0.038}$ $0.028_{\pm 0.000}$
MIWAE $0.011_{\pm 0.001}$ $0.827_{\pm 0.029}$ $0.036_{\pm 0.001}$
not-MIWAE $0.009_{\pm 0.000}$ $0.845_{\pm 0.033}$ $0.033_{\pm 0.001}$
EGC $0.006_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.062_{\pm 0.000}$
MaCoDE $0.009_{\pm 0.001}$ $0.864_{\pm 0.034}$ $0.037_{\pm 0.001}$
Table 13: Q3: Multiple imputation for each dataset under MCAR at 0.3 missingness. The means and standard errors of the mean across 5 datasets and 10 repeated experiments are reported. $\downarrow$ denotes lower is better.
Dataset abalone
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.005_{\pm 0.001}$ $0.971_{\pm 0.019}$ $0.030_{\pm 0.000}$
GAIN $0.011_{\pm 0.002}$ $0.729_{\pm 0.072}$ $0.030_{\pm 0.000}$
missMDA $0.015_{\pm 0.001}$ $0.571_{\pm 0.048}$ $0.032_{\pm 0.001}$
VAEAC $0.003_{\pm 0.001}$ $0.986_{\pm 0.014}$ $0.031_{\pm 0.000}$
MIWAE $0.005_{\pm 0.001}$ $0.971_{\pm 0.019}$ $0.033_{\pm 0.001}$
not-MIWAE $0.006_{\pm 0.001}$ $0.943_{\pm 0.023}$ $0.032_{\pm 0.001}$
EGC $0.004_{\pm 0.001}$ $0.986_{\pm 0.014}$ $0.035_{\pm 0.001}$
MaCoDE $0.005_{\pm 0.001}$ $0.957_{\pm 0.022}$ $0.033_{\pm 0.001}$
Dataset banknote
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.012_{\pm 0.002}$ $0.800_{\pm 0.033}$ $0.052_{\pm 0.000}$
GAIN $0.021_{\pm 0.004}$ $0.675_{\pm 0.084}$ $0.052_{\pm 0.000}$
missMDA $0.014_{\pm 0.001}$ $0.800_{\pm 0.033}$ $0.055_{\pm 0.000}$
VAEAC $0.009_{\pm 0.001}$ $0.925_{\pm 0.038}$ $0.053_{\pm 0.000}$
MIWAE $0.004_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.054_{\pm 0.000}$
not-MIWAE $0.004_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.054_{\pm 0.000}$
EGC $0.005_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.071_{\pm 0.003}$
MaCoDE $0.004_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.054_{\pm 0.000}$
Dataset breast
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.005_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.080_{\pm 0.000}$
GAIN $0.011_{\pm 0.001}$ $0.960_{\pm 0.016}$ $0.080_{\pm 0.000}$
missMDA $0.019_{\pm 0.001}$ $0.887_{\pm 0.015}$ $0.084_{\pm 0.000}$
VAEAC $0.006_{\pm 0.000}$ $0.997_{\pm 0.002}$ $0.081_{\pm 0.000}$
MIWAE $0.004_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.082_{\pm 0.000}$
not-MIWAE $0.005_{\pm 0.001}$ $0.993_{\pm 0.007}$ $0.081_{\pm 0.000}$
EGC $0.006_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.083_{\pm 0.000}$
MaCoDE $0.007_{\pm 0.000}$ $0.997_{\pm 0.003}$ $0.083_{\pm 0.000}$
Dataset redwine
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.015_{\pm 0.001}$ $0.836_{\pm 0.035}$ $0.048_{\pm 0.000}$
GAIN $0.020_{\pm 0.001}$ $0.627_{\pm 0.021}$ $0.048_{\pm 0.000}$
missMDA $0.020_{\pm 0.002}$ $0.709_{\pm 0.035}$ $0.053_{\pm 0.000}$
VAEAC $0.011_{\pm 0.001}$ $0.882_{\pm 0.019}$ $0.049_{\pm 0.000}$
MIWAE $0.008_{\pm 0.001}$ $0.945_{\pm 0.024}$ $0.053_{\pm 0.001}$
not-MIWAE $0.009_{\pm 0.001}$ $0.909_{\pm 0.030}$ $0.051_{\pm 0.000}$
EGC $0.007_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.067_{\pm 0.001}$
MaCoDE $0.008_{\pm 0.000}$ $0.980_{\pm 0.013}$ $0.054_{\pm 0.001}$
Dataset whitewine
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.008_{\pm 0.001}$ $0.773_{\pm 0.034}$ $0.028_{\pm 0.000}$
GAIN $0.025_{\pm 0.003}$ $0.500_{\pm 0.053}$ $0.028_{\pm 0.000}$
missMDA $0.011_{\pm 0.000}$ $0.718_{\pm 0.032}$ $0.033_{\pm 0.001}$
VAEAC $0.008_{\pm 0.001}$ $0.827_{\pm 0.029}$ $0.029_{\pm 0.000}$
MIWAE $0.007_{\pm 0.001}$ $0.891_{\pm 0.030}$ $0.033_{\pm 0.001}$
not-MIWAE $0.006_{\pm 0.001}$ $0.945_{\pm 0.024}$ $0.031_{\pm 0.000}$
EGC $0.006_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.058_{\pm 0.001}$
MaCoDE $0.007_{\pm 0.003}$ $0.882_{\pm 0.075}$ $0.033_{\pm 0.003}$
Table 14: Q3: Multiple imputation for each dataset under MAR at 0.3 missingness. The means and standard errors of the mean across 5 datasets and 10 repeated experiments are reported. $\downarrow$ denotes lower is better. $\downarrow$ denotes lower is better.
Dataset abalone
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.007_{\pm 0.001}$ $0.929_{\pm 0.024}$ $0.030_{\pm 0.000}$
GAIN $0.015_{\pm 0.002}$ $0.629_{\pm 0.061}$ $0.030_{\pm 0.000}$
missMDA $0.018_{\pm 0.001}$ $0.471_{\pm 0.057}$ $0.034_{\pm 0.001}$
VAEAC $0.005_{\pm 0.001}$ $0.957_{\pm 0.022}$ $0.030_{\pm 0.000}$
MIWAE $0.008_{\pm 0.001}$ $0.900_{\pm 0.043}$ $0.034_{\pm 0.001}$
not-MIWAE $0.010_{\pm 0.002}$ $0.886_{\pm 0.060}$ $0.035_{\pm 0.003}$
EGC $0.005_{\pm 0.000}$ $0.986_{\pm 0.014}$ $0.039_{\pm 0.001}$
MaCoDE $0.007_{\pm 0.001}$ $0.929_{\pm 0.024}$ $0.035_{\pm 0.001}$
Dataset banknote
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.017_{\pm 0.001}$ $0.775_{\pm 0.025}$ $0.053_{\pm 0.000}$
GAIN $0.032_{\pm 0.005}$ $0.625_{\pm 0.077}$ $0.053_{\pm 0.000}$
missMDA $0.019_{\pm 0.001}$ $0.725_{\pm 0.045}$ $0.056_{\pm 0.000}$
VAEAC $0.010_{\pm 0.001}$ $0.975_{\pm 0.025}$ $0.053_{\pm 0.000}$
MIWAE $0.005_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.055_{\pm 0.000}$
not-MIWAE $0.005_{\pm 0.003}$ $1.000_{\pm 0.000}$ $0.055_{\pm 0.000}$
EGC $0.007_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.072_{\pm 0.002}$
MaCoDE $0.006_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.054_{\pm 0.000}$
Dataset breast
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.008_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.080_{\pm 0.000}$
GAIN $0.014_{\pm 0.001}$ $0.960_{\pm 0.011}$ $0.080_{\pm 0.000}$
missMDA $0.026_{\pm 0.001}$ $0.853_{\pm 0.015}$ $0.086_{\pm 0.000}$
VAEAC $0.009_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.082_{\pm 0.000}$
MIWAE $0.006_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.084_{\pm 0.000}$
not-MIWAE $0.007_{\pm 0.002}$ $0.990_{\pm 0.016}$ $0.082_{\pm 0.000}$
EGC $0.007_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.084_{\pm 0.000}$
MaCoDE $0.009_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.084_{\pm 0.000}$
Dataset redwine
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.021_{\pm 0.001}$ $0.673_{\pm 0.034}$ $0.048_{\pm 0.000}$
GAIN $0.026_{\pm 0.002}$ $0.473_{\pm 0.057}$ $0.048_{\pm 0.000}$
missMDA $0.028_{\pm 0.000}$ $0.600_{\pm 0.028}$ $0.055_{\pm 0.000}$
VAEAC $0.016_{\pm 0.001}$ $0.782_{\pm 0.036}$ $0.050_{\pm 0.000}$
MIWAE $0.012_{\pm 0.001}$ $0.909_{\pm 0.019}$ $0.055_{\pm 0.001}$
not-MIWAE $0.013_{\pm 0.001}$ $0.855_{\pm 0.047}$ $0.053_{\pm 0.001}$
EGC $0.009_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.072_{\pm 0.001}$
MaCoDE $0.011_{\pm 0.001}$ $0.990_{\pm 0.010}$ $0.055_{\pm 0.001}$
Dataset whitewine
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.012_{\pm 0.001}$ $0.673_{\pm 0.039}$ $0.028_{\pm 0.000}$
GAIN $0.029_{\pm 0.004}$ $0.427_{\pm 0.045}$ $0.028_{\pm 0.000}$
missMDA $0.015_{\pm 0.000}$ $0.582_{\pm 0.034}$ $0.035_{\pm 0.001}$
VAEAC $0.012_{\pm 0.001}$ $0.691_{\pm 0.024}$ $0.029_{\pm 0.000}$
MIWAE $0.010_{\pm 0.001}$ $0.873_{\pm 0.024}$ $0.036_{\pm 0.001}$
not-MIWAE $0.008_{\pm 0.001}$ $0.891_{\pm 0.072}$ $0.033_{\pm 0.002}$
EGC $0.008_{\pm 0.000}$ $0.991_{\pm 0.009}$ $0.063_{\pm 0.001}$
MaCoDE $0.009_{\pm 0.001}$ $0.827_{\pm 0.037}$ $0.034_{\pm 0.001}$
Table 15: Q3: Multiple imputation for each dataset under MNARL at 0.3 missingness. The means and standard errors of the mean across 5 datasets and 10 repeated experiments are reported. $\downarrow$ denotes lower is better.
Dataset abalone
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.003_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.030_{\pm 0.000}$
GAIN $0.008_{\pm 0.001}$ $0.886_{\pm 0.036}$ $0.031_{\pm 0.001}$
missMDA $0.009_{\pm 0.000}$ $0.800_{\pm 0.032}$ $0.033_{\pm 0.000}$
VAEAC $0.002_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.030_{\pm 0.000}$
MIWAE $0.005_{\pm 0.001}$ $0.957_{\pm 0.022}$ $0.032_{\pm 0.001}$
not-MIWAE $0.007_{\pm 0.001}$ $0.943_{\pm 0.023}$ $0.035_{\pm 0.001}$
EGC $0.003_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.038_{\pm 0.001}$
MaCoDE $0.007_{\pm 0.001}$ $0.886_{\pm 0.019}$ $0.033_{\pm 0.001}$
Dataset banknote
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.008_{\pm 0.001}$ $0.875_{\pm 0.042}$ $0.052_{\pm 0.000}$
GAIN $0.013_{\pm 0.002}$ $0.825_{\pm 0.053}$ $0.052_{\pm 0.000}$
missMDA $0.009_{\pm 0.001}$ $0.775_{\pm 0.045}$ $0.054_{\pm 0.000}$
VAEAC $0.005_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.053_{\pm 0.000}$
MIWAE $0.003_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.054_{\pm 0.000}$
not-MIWAE $0.003_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.054_{\pm 0.000}$
EGC $0.004_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.066_{\pm 0.001}$
MaCoDE $0.003_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.053_{\pm 0.000}$
Dataset breast
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.004_{\pm 0.000}$ $0.997_{\pm 0.003}$ $0.080_{\pm 0.000}$
GAIN $0.007_{\pm 0.000}$ $0.997_{\pm 0.003}$ $0.080_{\pm 0.000}$
missMDA $0.017_{\pm 0.001}$ $0.927_{\pm 0.012}$ $0.084_{\pm 0.000}$
VAEAC $0.006_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.082_{\pm 0.000}$
MIWAE $0.004_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.082_{\pm 0.000}$
not-MIWAE $0.004_{\pm 0.000}$ $0.997_{\pm 0.003}$ $0.081_{\pm 0.000}$
EGC $0.004_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.082_{\pm 0.000}$
MaCoDE $0.005_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.083_{\pm 0.000}$
Dataset redwine
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.010_{\pm 0.001}$ $0.873_{\pm 0.028}$ $0.048_{\pm 0.000}$
GAIN $0.011_{\pm 0.001}$ $0.827_{\pm 0.032}$ $0.048_{\pm 0.000}$
missMDA $0.016_{\pm 0.002}$ $0.755_{\pm 0.024}$ $0.052_{\pm 0.000}$
VAEAC $0.009_{\pm 0.001}$ $0.900_{\pm 0.025}$ $0.050_{\pm 0.000}$
MIWAE $0.006_{\pm 0.001}$ $0.955_{\pm 0.020}$ $0.052_{\pm 0.000}$
not-MIWAE $0.006_{\pm 0.001}$ $0.973_{\pm 0.014}$ $0.050_{\pm 0.000}$
EGC $0.005_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.063_{\pm 0.001}$
MaCoDE $0.006_{\pm 0.001}$ $0.982_{\pm 0.018}$ $0.053_{\pm 0.000}$
Dataset whitewine
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.008_{\pm 0.000}$ $0.773_{\pm 0.020}$ $0.028_{\pm 0.000}$
GAIN $0.012_{\pm 0.002}$ $0.736_{\pm 0.029}$ $0.028_{\pm 0.000}$
missMDA $0.010_{\pm 0.000}$ $0.809_{\pm 0.021}$ $0.033_{\pm 0.001}$
VAEAC $0.009_{\pm 0.001}$ $0.736_{\pm 0.029}$ $0.028_{\pm 0.000}$
MIWAE $0.006_{\pm 0.001}$ $0.964_{\pm 0.020}$ $0.033_{\pm 0.001}$
not-MIWAE $0.006_{\pm 0.000}$ $0.918_{\pm 0.025}$ $0.031_{\pm 0.001}$
EGC $0.004_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.055_{\pm 0.001}$
MaCoDE $0.006_{\pm 0.001}$ $0.855_{\pm 0.031}$ $0.032_{\pm 0.001}$
Table 16: Q3: Multiple imputation for each dataset under MNARQ at 0.3 missingness. The means and standard errors of the mean across 5 datasets and 10 repeated experiments are reported. $\downarrow$ denotes lower is better.